### v1

# The Reward Model Selection Crisis in Personalized Alignment

Abstract
Personalized alignment from preference data has focused primarily on improving reward model (RM) accuracy, with the implicit assumption that better preference ranking translates to better personalized behavior. However, in deployment, computational constraints necessitate inference-time adaptation via reward-guided decoding (RGD) rather than per-user policy fine-tuning. This creates a critical but overlooked requirement: reward models must not only rank preferences accurately but also effectively guide token-level generation decisions. We demonstrate that standard RM accuracy fails catastrophically as a selection criterion for deployment-ready personalized alignment. Through systematic evaluation across three datasets, we introduce policy accuracy—a metric quantifying whether RGD scoring functions correctly discriminate between preferred and dispreferred responses—and show that RM accuracy correlates only weakly with this discrimination ability (Kendall’s $\tau$ = 0.08–0.31). More critically, we introduce Pref-LaMP, the first personalized alignment benchmark with ground-truth user completions, enabling direct behavioral evaluation without circular reward-based metrics. On Pref-LaMP, we expose a complete decoupling between discrimination and generation: methods with 20-point RM accuracy differences produce almost identical output quality, and even methods achieving high discrimination fail to generate behaviorally aligned responses. Furthermore, simple in-context learning (ICL) dominates all reward-guided methods for models $\geq$3B parameters, achieving 3–5 point ROUGE-1 gains over the best reward method at 7B scale. These findings reveal that the field has been optimizing for proxy metrics that don’t predict deployment performance, and that current personalized alignment methods fundamentally fail to operationalize preferences into behavioral adaptation under realistic deployment constraints.
The Reward Model Selection Crisis in Personalized Alignment
Fady Rezk1,2, Yuangang Pan2, Chuan-Sheng Foo2, Xun Xu2, Nancy Chen2, Henry Gouk1, Timothy Hospedales1,3 1 School of Informatics, University of Edinburgh, UK 2 Agency for Science, Technology and Research (A*STAR), Singapore 3 Samsung AI Research Center, Cambridge, UK
1 Introduction
Recent advances in aligning large language models (LLMs) with human preferences have primarily focused on learning from aggregated feedback across diverse user populations (Rafailov et al., 2023; Ouyang et al., 2022). However, preferences are inherently pluralistic—varying across individuals, communities, and contexts (Santurkar et al., 2023; Sorensen et al., 2024). This reality motivates personalized alignment: adapting model behavior to heterogeneous, sometimes conflicting, user preferences rather than collapsing them into a single consensus objective.
Current personalized alignment research has converged on a common paradigm: collect user-specific preference data (pairwise comparisons, ratings, or implicit signals), train personalized reward models to capture individual preferences, and assume that better reward models naturally translates to better policies (Bose et al., 2025; Chen et al., 2025a; Shenfeld et al., 2025; Li et al., 2024; Poddar et al., 2024). The last assumption is known to break as reasoned by Goodhart’s law El-Mhamdi and Hoang (2024). Unlike standard RLHF, there are no downstream benchmark that measures policy performance such as MMLU Hendrycks et al. (2021) and GSM8k Cobbe et al. (2021).
1.1 The Deployment Reality and End-to-End Perspective
Per-user policy fine-tuning using reward model artefacts via RL is computationally infeasible at scale. For 1000 user groups, RL-based personalization requires dynamic adapter management, RL instability mitigation, and orders of magnitude more compute than inference-time alternatives. The only scalable path is inference-time adaptation through reward-guided decoding (RGD) Khanov et al. (2024), maintaining a single base policy while using personalized rewards to guide generation. Another option is or Best-of-N sampling Ichihara et al. (2025) but the high latency of BoN makes it unfit for personalization text generation.
This deployment reality demands we adopt an end-to-end behavioral perspective: personalized alignment is not merely reward modeling, but the complete process from preference data to actual generation behavior. We propose a key principle:
A personalized alignment method must specify not only how preferences are modeled, but how they are operationalized into behavioral adaptation. Papers proposing reward models are responsible for demonstrating that improved RM accuracy translates to improved generation behavior.
Current evaluation ignores this responsibility, treating reward modeling and policy adaptation as isolated steps. This obscures whether methods actually achieve their objective: making models generate responses aligned with user preferences.
1.2 Our Investigation
We adopt an end-to-end perspective, studying the complete chain from preference modeling to generation behavior. We ask three questions: (1) Does RM accuracy predict discrimination under RGD? (2) Does discrimination predict generation quality? (3) How do reward methods compare to simpler baselines?
To answer these, we introduce policy accuracy—measuring whether RGD scoring assigns higher scores to preferred responses—and Pref-LaMP, a benchmark with ground-truth user completions enabling direct behavioral evaluation.
Our findings reveal a fundamental selection crisis: methods with 20-point RM accuracy differences produce identical generation quality, and correlations between RM accuracy and downstream performance are negligibly weak (Kendall’s $\tau$ = 0.08-0.31). Practitioners cannot reliably choose deployment-ready methods because standard metrics don’t predict actual performance.
1.3 Key Findings
Finding 1: RM accuracy doesn’t predict discrimination (Kendall’s $\tau$ = 0.08–0.31). Methods with 20-point RM accuracy differences achieve nearly identical policy performance.
Finding 2: Discrimination doesn’t predict generation. On Pref-LaMP, methods cluster tightly in generation quality despite dramatic variance in RM and policy accuracy.
Finding 3: ICL dominates at scale. At 3B-7B parameters, ICL-RAG outperforms all reward methods by 3-5 ROUGE-1 points.
Implications: For practitioners, use ICL-RAG at scale. For researchers, RM accuracy is broken—evaluate on ground-truth completions. The field needs behavioral benchmarks like Pref-LaMP, analogous to GSM8K/MMLU for general RLHF.
1.4 Contributions
We (1) formalize the RM selection problem and introduce policy accuracy for discrimination measurement, (2) contribute Pref-LaMP—the first benchmark with ground-truth user completions, (3) demonstrate RM accuracy fails as a selection criterion across three datasets and four scales, and (4) provide actionable recommendations for practitioners and researchers.
2 Related Work
2.1 Personalized Alignment
Recent work has focused on learning user-specific reward models or policies for alignment under limited supervision (e.g., PAL, PReF, LoRE, P-DPO, VPL) (Chen et al., 2025a; Poddar et al., 2024; Bose et al., 2025; Shenfeld et al., 2025; Li et al., 2024). These approaches largely target reward-modeling accuracy or relative preference ranking as proxies for personalization quality Chen et al. (2025a); Bose et al. (2025); Shenfeld et al. (2025). However, such metrics often fail to capture whether a policy’s generated responses truly reflect an individual user’s preferences, especially under resource-constrained settings or whether it is an unintended overoptimization (Pan et al., 2022).
2.2 Multi-Objective Alignment
Multi-objective alignment (MOA) addresses the challenge of optimizing language models across multiple known and predefined reward dimensions simultaneously. Unlike personalized alignment, where the goal is to learn individual user preferences under limited supervision, MOA assumes access to distinct reward models for each objective dimension (e.g., helpfulness, harmlessness, factuality) and focuses on finding optimal policy trade-offs among these objectives. Prior work has explored weighted reward optimization (Zhou et al., 2024), model merging (Jang et al., 2024; Rame et al., 2023), auxiliary correction models (Ji et al., 2024; Yang et al., 2024a), test-time reward-guided decoding Chen et al. (2025b) among other methods (Yang et al., 2024b). Despite the terminological overlap with ”personalization” in some methods, the key distinction is that MOA operates over known objective dimensions with pre-trained reward models, whereas personalized alignment learns latent user-specific preferences from sparse behavioral data across dataset specific dimensions (e.g., style, length, political values, rhetoric, culture sensitivities, etc).
2.3 Evaluation Challenges in RLHF
Existing evaluation practices in RLHF and personalized alignment rely on proxy metrics such as reward-model scores which are susceptible to reward hacking and circularity (Tien et al., 2023; Pan et al., 2022). These methods assess optimization success rather than behavioral fidelity Wen et al. (2025); Gao et al. (2023); El-Mhamdi and Hoang (2024). In contrast, our work introduces a framework for direct behavioral evaluation, measuring whether generated responses match user-provided completions (Section 5). Extended discussion of related evaluation pathologies appears in Appendix A.
Our work addresses these limitations by introducing direct behavioral evaluation on ground-truth user completions, measuring actual generation quality rather than relying on proxy metrics. Extended discussion of evaluation pathologies appears in Appendix A.
2.4 Inference-Time Alignment
Reward-guided decoding (ARGS) Khanov et al. (2024) and Best-of-N sampling Ichihara et al. (2025) enable policy steering without fine-tuning, making them computationally attractive for personalization. Recent work has explored their effectiveness Wu (2025), but evaluation remains limited to reward-based metrics. Our work is the first to systematically evaluate these methods for personalized alignment with ground-truth behavioral assessment, revealing fundamental limitations in their ability to operationalize user preferences.
3 Preliminaries and Problem Setup
3.1 Personalized Preference Learning
Consider a preference dataset $\mathcal{D}=\{(u_{i},x_{i},y_{i}^{(c)},y_{i}^{(r)})\}_{i=1}^{N}$, where $u_{i}\in U=\{1..K\}$ is a user identifier, $x_{i}$ is a prompt, and $y_{i}^{(c)},y_{i}^{(r)}$ are chosen and rejected completions. We split users into training $\mathcal{U}_{\mathrm{train}}$ (learning shared structure) and adaptation $\mathcal{U}_{\mathrm{adapt}}$ (few-shot personalization at deployment). For user $k$, define $\mathcal{D}_{k}=\{(x_{i},y_{i}^{(c)},y_{i}^{(r)}):u_{i}=k\}$.
Personalized Reward Modeling. The reward model conditions on user identity: $r_{\theta,z_{k}}(y\mid x)$, decomposing into shared parameters $\theta$ (general preference structure) and user-specific parameters $z_{k}$ (individual preferences). Training on $\mathcal{U}_{\mathrm{train}}$ learns both $\theta$ and $\{z_{k}\}_{k\in\mathcal{U}_{\mathrm{train}}}$. At deployment, for new user $k\in\mathcal{U}_{\mathrm{adapt}}$ with few-shot data $\mathcal{D}_{k}^{\mathrm{few-shot}}$, we adapt:
$z_{k}=\mathcal{A}_{R}\big(\mathcal{D}_{k}^{\mathrm{few-shot}};\theta\big)$
where $\mathcal{A}_{R}$ is the adaptation algorithm.
3.2 Deployment via Reward-Guided Decoding
Given computational constraints prohibiting per-user policy fine-tuning, we deploy personalized alignment through inference-time guidance. We consider two methods:
Reward-Guided Decoding (ARGS). At each generation step $t$, ARGS scores the top-$k$ tokens $V^{(k)}_{t}$ as:
$\text{score}(v\mid x,y_{<t};z_{k},\lambda)$ $=\log\pi(v\mid x,y_{<t})$ (1)
$\quad+\lambda\cdot r_{\theta,z_{k}}(v\mid x,y_{<t})$
selecting $y_{t}=\operatorname*{arg\,max}_{v\in V^{(k)}_{t}}\text{score}(v)$. This balances base model fluency ($\log\pi$) with personalized alignment ($r_{\theta,z_{k}}$).
3.3 The Evaluation Gap
Standard practice evaluates personalization methods by reward model accuracy:
Definition 1 (Reward Model Ranking Accuracy).
For user $k$ with evaluation set $\mathcal{D}_{k}^{\text{eval}}=\{(x_{i},y_{i}^{(c)},y_{i}^{(r)})\}$:
$\frac{1}{|\mathcal{D}_{k}^{\text{eval}}|}\sum_{\mathcal{D}_{k}^{\text{eval}}}\mathbb{I}\Big[r_{\theta,z_{k}}(y_{i}^{(c)}\mid x_{i})>r_{\theta,z_{k}}(y_{i}^{(r)}\mid x_{i})\Big]$ (2)
This measures pairwise ranking on complete responses. However, deployment via RGD requires token-level scoring. A reward model achieving high sequence-level ranking accuracy may provide poor token-level guidance, yet standard evaluation cannot detect this mismatch.
4 Policy Accuracy: Measuring Discrimination Under RGD
To bridge the evaluation gap, we introduce a metric quantifying whether the RGD scoring function—not just the reward model in isolation—can discriminate preferred from dispreferred responses.
Definition 2 (Policy Ranking Accuracy).
Let $s_{\pi}:\mathcal{Y}\times\mathcal{X}\to\mathbb{R}$ be the policy scoring function used at generation time. The policy accuracy for user $k$ is:
$\frac{1}{|\mathcal{D}_{k}^{\text{eval}}|}\sum_{\mathcal{D}_{k}^{\text{eval}}}\mathbb{I}\Big[s_{\pi}(y_{i}^{(c)}\mid x_{i})>s_{\pi}(y_{i}^{(r)}\mid x_{i})\Big]$ (3)
We instantiate $s_{\pi}$ with three scoring functions, each revealing different aspects of the personalization pipeline:
Prior Policy. The base policy’s length-normalized log-likelihood, quantifying zero-shot discrimination ability:
$\frac{1}{|y|}\sum_{t=1}^{|y|}\log\pi(y_{t}\mid x,y_{<t})$ (4)
Global RM Posterior. RGD scoring with a non-personalized reward model:
$\sum_{t=1}^{|y|}\Big[\log\pi(y_{t}\mid x,y_{<t})+\lambda\cdot r_{\theta}(y_{t}\mid x,y_{<t})\Big]$ (5)
Personalized Posterior. RGD scoring with personalized reward model:
$\sum_{t=1}^{|y|}\Big[\log\pi(y_{t}\mid x,y_{<t})$ $\quad+\lambda\cdot r_{\theta,z_{k}}(y_{t}\mid x,y_{<t})\Big]$ (6)
Comparing these accuracies reveals: (1) whether reward guidance improves discrimination over the base policy, and (2) whether personalization provides gains beyond global reward models.
Critical Limitation. Policy accuracy measures discrimination ability—whether the scoring function prefers chosen responses—not generation quality. High policy accuracy provides no guarantee the policy will generate outputs that behaviorally align with user preferences. A method might correctly rank static responses while producing generations that differ substantially from what users would actually write. This motivates our ground-truth behavioral evaluation in Section 5.
5 Pref-LaMP: A Benchmark for Direct Behavioral Evaluation
To enable direct measurement of behavioral alignment without circular reward-based metrics, we introduce Pref-LaMP—a personalized alignment benchmark providing both pairwise preferences and ground-truth user-authored completions.
5.1 Dataset Construction
Pref-LaMP derives from LaMP-5 Salemi et al. (2024), pairing researchers’ abstracts with their titles. Both are author-written, capturing individual style. We construct preferences via hard negative mining: (1) encode abstracts with Qwen3-Embedding-0.6B, (2) retrieve top-$k$ similar abstracts from other users, (3) sample one retrieved abstract and use its title as $y^{(r)}$, (4) use original title as $y^{(c)}$. This ensures rejections are topically relevant but stylistically different. Human-written negatives avoid shortcut learning—initial LLM-generated rejections let linear probes detect generation artifacts rather than preference signals.
Pref-LaMP is the first benchmark enabling non-circular evaluation through user-authored completions, measurable via ROUGE and BERTScore.
5.2 Behavioral Alignment Metric
Definition 3 (Behavioral Alignment).
Let $\mathcal{G}:\mathcal{X}\to\mathcal{Y}$ be a generation operator and $\mathcal{S}:\mathcal{Y}\times\mathcal{Y}\to[0,1]$ be a similarity measure. For user $k$ with test set $\mathcal{P}_{k}=\{(x_{j},y_{j}^{\text{GT}})\}_{j=1}^{M_{k}}$:
$\mathcal{A}_{\mathcal{S}}(\mathcal{G},k)=\frac{1}{|\mathcal{P}_{k}|}\sum_{(x_{j},y_{j}^{\text{GT}})\in\mathcal{P}_{k}}\mathcal{S}\Big(\mathcal{G}(x_{j}),y_{j}^{\text{GT}}\Big)$ (7)
We instantiate $\mathcal{G}$ with ARGS decoding (Eq. 1), and $\mathcal{S}$ with ROUGE-1 (lexical overlap), ROUGE-L (longest common subsequence), and BERTScore-F1 (semantic similarity). This measures whether generated outputs match what users actually write, providing ground-truth assessment of behavioral alignment.
6 Experimental Setup
6.1 Datasets and Models
TLDR Stiennon et al. (2020): Binary stylistic preferences, 10 training users (2,097 prefs/user), 31 adaptation users (100 prefs/user). PRISM Kirk et al. (2024): Pluralistic preferences, 1,232 training users (22.1 prefs/user), 139 adaptation users (14.5 prefs/user). Pref-LaMP (ours): User-authored completions, 485 training users (48.8 prefs/user), 126 adaptation users (49.2 prefs/user).
6.2 Models and Methods
Dataset Reward Model Base Policies
TLDR SmolLM2-180M 180M, 360M, 1.7B
PRISM Qwen2.5-0.5B 0.5B, 1.5B, 3B, 7B
Pref-LaMP Qwen2.5-0.5B 0.5B, 1.5B, 3B, 7B
Table 1: Model configurations.
All reward models use LoRA rank 8, trained on $\mathcal{U}_{\text{train}}$ to learn shared $\theta$ and user-specific $\{z_{k}\}$. We evaluate six personalization methods: LoRE Bose et al. (2025) (learn reward bases and user-specific convex combination), LoRE-Alt (same as LoRE but alternates between bases and user specific parameter gradient steps), PReF Shenfeld et al. (2025) (collaborative filtering), PAL Chen et al. (2025a), VPL Poddar et al. (2024), MPU/MPU-Avg (a simple baselines of per-user MLPs), and P-DPO Li et al. (2024) (personalized direct preference optimization). Baselines include Global-RM (non-personalized Bradley-Terry using last token embedding as input to RM decoder), Global-RM-V2 (a sequence reward is average reward for all tokens), GenARM Xu et al. (2025) (autoregressive RM for token-level guidance), zero-shot generation, ICL (random demonstrations), and ICL-RAG (retrieved demonstrations).
6.3 Evaluation Protocol
We measure: (1) RM accuracy on adaptation/deployment users held-out preferences, (2) policy accuracy using prior/global/personalized scoring functions, (3) generation quality on Pref-LaMP via ROUGE-1/L and BERTScore against ground-truth, and (4) win rates where each method’s RM judges its own outputs versus zero-shot baseline.
7 Results Part 1: Discrimination Does Not Follow From Ranking
We first investigate whether reward model accuracy predicts discrimination ability under reward-guided decoding—i.e., whether methods that rank preferences accurately can also guide generation appropriately.
7.1 TLDR: Weak Correlation on Simple Data
Policy Accuracy
Method RM Accuracy 135M 360M 1.7B
LoRE 82.56 $\pm$ 1.12 81.09 $\pm$ 0.76 81.09 $\pm$ 0.89 81.01 $\pm$ 1.10
MPU 76.76 $\pm$ 0.26 49.96 $\pm$ 3.24 54.36 $\pm$ 7.36 55.12 $\pm$ 8.07
MPU-Avg 77.83 $\pm$ 0.19 52.91 $\pm$ 4.01 52.96 $\pm$ 3.97 52.94 $\pm$ 3.98
P-DPO - - - -
PAL 87.00 $\pm$ 0.05 - - -
PReF 65.27 $\pm$ 8.15 46.68 $\pm$ 7.18 49.56 $\pm$ 11.77 49.65 $\pm$ 11.95
VPL 68.67 $\pm$ 1.94 59.78 $\pm$ 4.34 60.40 $\pm$ 4.82 60.67 $\pm$ 4.92
Global 73.67 $\pm$ 2.45 58.80 $\pm$ 4.79 59.58 $\pm$ 1.55 60.34 $\pm$ 1.80
Global-V2 38.66 $\pm$ 1.40 53.05 $\pm$ 0.02 61.26 $\pm$ 0.07 62.54 $\pm$ 0.04
GenARM 53.67 $\pm$ 3.15 53.06 $\pm$ 0.02 61.56 $\pm$ 0.03 62.59 $\pm$ 0.02
Prior - 62.00 $\pm$ 0.00 61.57 $\pm$ 0.00 64.60 $\pm$ 0.00
Pearson $r$ - 0.41 0.15 0.11
Spearman $\rho$ - 0.36 0.09 0.09
Kendall $\tau$ - 0.31 0.08 0.09
Table 2: Policy accuracy and reward model performance on TLDR + SmolLM. Base reward model is SmolLM2-135M
On TLDR’s binary preferences, most personalization methods (MPU, MPU-Avg, PReF) achieve policy accuracies below the prior baseline (62-65%), particularly at smaller scales. RM accuracy correlates weakly with policy accuracy (Kendall’s $\tau$ = 0.08–0.31), degrading as scale increases (Pearson $r$: 0.41 $\to$ 0.11 from 180M to 1.7B). PAL achieves highest RM accuracy (87.0%) but doesn’t support RGD, while LoRE—with lower RM accuracy (82.6%)—achieves superior policy accuracy ($\sim$81%). LoRE substantially outperforms all baselines with 16-20 point gains over prior policy, demonstrating genuine effectiveness and validating our evaluation framework.
7.2 PRISM: Personalization Fails on Pluralistic Data
LoRE LoRE-Alt MPU MPU-Avg PAL PReF VPL Global RM Global RM-v2
RM Accuracy $65.95$ $\pm 4.57$ $67.15$ $\pm 0.42$ $52.76$ $\pm 5.40$ $60.43$ $\pm 1.25$ $70.74$ $\pm 0.42$ $74.34$ $\pm 2.20$ $68.35$ $\pm 2.59$ 77.94 $\pm 3.55$ $62.11$ $\pm 4.15$
Table 3: Reward model accuracies of various algorithms using Qwen2.5-0.5B backbone on PRISM dataset. Global reward model V1 outperforms all other methods.
Figure 1: RM vs Policy Accuracy Correlation on PRISM across model scales. Correlations remain consistently weak: Pearson $r$ ranges from 0.30–0.48, Spearman $\rho$ from 0.25–0.43, and Kendall $\tau$ from 0.17–0.29. While correlations slightly strengthen with scale, they remain far below what would be needed for RM accuracy to reliably predict policy performance. Notably, methods cluster vertically (similar RM accuracy, varying policy performance), demonstrating that high RM accuracy does not guarantee better policy understanding.
Results are summarized in Table 3 and Figure 1. With reference to RM accuracies in Table 3, all personalized methods underperform Global RM (77.9% vs. 52.8-74.3%), with even the best (PReF, 74.3%) falling 3.6 points short. Meanwhile, RM-policy correlation, as shown in Fig 1, remains weak (Kendall’s $\tau$ = 0.17-0.29), slightly strengthening with scale ($r$: 0.31 $\to$ 0.48). Critically, VPL achieves highest policy accuracy (63.8-66.4%) despite 9.5 points lower RM accuracy than Global RM—a complete ranking inversion. Methods show minimal scaling gains, remaining in narrow bands (LoRE: 56.8-57.6%, VPL: 63.8-66.4%), suggesting limitations in learning from sparse pluralistic.
Unlike TLDR/SmolLM2 where correlations degraded with scale ($r$ = 0.41→0.11), PRISM/Qwen2.5 shows strengthening correlations ($r$ = 0.31→0.48). Whether this reflects dataset differences, model architecture, or their interaction remains unclear. Regardless, even at 7B, correlations remain too weak for reliable prediction.
Reconciling with Prior Claims. LoRE’s released code fails to reproduce claimed gains over Global-RM [^1] https://github.com/facebookresearch/LoRe/issues/1; our reimplementation of their linear probe setup yields almost identical accuracy (us: 55.59% vs. theirs: 55.55%) validating our imlpementation. PReF was evaluated on PERSONA, not PRISM which they admittedly fail on (Appendix D). PAL and VPL used different datasets (TLDR, non-overlapping tasks).
Implication: RM accuracy cannot guide deployment—methods with 10+ point gaps perform identically or invert in ranking. This implies that RM accuracy does not measure behaviorally-relevant properties for policy guidance. Reward models can rank preferences accurately while failing to provide useful token-level guidance.
Recommendation. Evaluations must assess downstream policy understanding across multiple datasets and scales—not just RM accuracy. The RM-policy disconnect demands new metrics measuring reward models’ suitability for guiding generation, rather than pairwise ranking accuracy alone.
8 Results Part 2: Discrimination Does Not Imply Generation
Given the weak correlation between RM accuracy and policy discrimination ability under RGD, we now ask: even when methods achieve high policy accuracy—demonstrably preferring chosen over rejected responses—do they generate outputs that behaviorally align with user preferences?
8.1 Pref-LaMP: Complete Decoupling Revealed
LoRE LoRE-Alt MPU MPU-Avg PAL PReF VPL Vanilla Vanilla-v2 GenARM
RM Acc $65.60$ $\pm 7.89$ $84.96$ $\pm 0.48$ $65.26$ $\pm 1.03$ $67.30$ $\pm 0.30$ $53.77$ $\pm 0.28$ $51.46$ $\pm 3.84$ $43.63$ $\pm 3.77$ $84.96$ $\pm 0.13$ $84.69$ $\pm 0.07$ $83.89$ $\pm 0.60$
Table 4: Reward model accuracies using Qwen2.5-0.5B on Pref-LaMP. Vanilla RMs again outperform personalization methods.
Figure 2: Generation quality (ROUGE-1, ROUGE-L, BERTScore-F1) under RGD across model scales. At 0.5B-1.5B, personalized RMs marginally improve over zero-shot; at 3B-7B, ICL baselines dominate all reward-guided methods.
Global RMs dominate (84.7-85.0%), with GenARM (83.9%) and LoRE-Alt close. Most personalization methods fail (VPL: 43.6%, PReF: 51.5%). Figure 2 reveals three critical failures:
(1) Architecture provides no advantage. GenARM, designed for token-level guidance, performs comparably to vanilla methods despite similar RM accuracy.
(2) RM accuracy decouples from generation. Best reward methods (LoRE, Vanilla RM, MPU-Avg) achieve identical ROUGE-1 (45.9-46.0) and BERTScore (70.5-70.8) at 7B despite 20-point RM accuracy gaps. High RM accuracy guarantees nothing about generation quality.
(3) Reward guidance fails at scale. At 0.5-1.5B, reward methods improve over zero-shot. At 3-7B, ICL-RAG dominates, achieving 48.9 ROUGE-1 and 72.2 BERTScore at 7B—3.0 and 1.5 point gaps over best reward methods. RGD becomes counterproductive as base capability increases.
Method 0.5B 1.5B 3B 7B
GenARM 100.0% 100.0% 100.0% 99.8%
LoRE 73.7% 72.9% 73.8% 73.7%
LoRE-Alt 97.0% 96.0% 96.3% 96.3%
MPU 61.9% 61.8% 61.5% 61.5%
MPU-Avg 69.8% 69.3% 70.1% 69.9%
PReF 54.9% 54.8% 56.1% 55.5%
VPL 51.1% 50.1% 50.5% 48.4%
Vanilla RM 99.9% 99.9% 99.9% 99.9%
Vanilla RM-v2 98.4% 97.2% 89.4% 84.4%
Table 5: Win rate: fraction of examples where each RM judges its own RGD output as better than zero-shot. High win rates reveal severe reward hacking—methods claim near-perfect improvement despite ground-truth metrics showing minimal or negative gains.
Win rates expose reward hacking: GenARM claims 100% improvement over zero-shot despite ROUGE-1 being marginally worse. Vanilla RM claims 99.9% superiority while producing outputs identical to methods 20 points lower in RM accuracy. Circular evaluation actively misleads—methods optimize their scoring functions then declare success while actual alignment remains unmeasured.
Figure 3: ROUGE-1 vs. number of in-context examples. ICL-RAG consistently outperforms random ICL, with performance scaling with both model size and context length.
ICL-RAG improves steadily with demonstrations and scale, reaching $\sim$49 ROUGE-1 at 7B with 5 shots. Larger models show no saturation, effectively leveraging context without reward guidance. This reveals RGD’s failure is not inherent to personalization but to its operationalization mechanism.
8.2 No Metric Predicts Generation Quality
Scale RM Acc Policy Acc Win Rate
Kendall’s $\tau$ with ROUGE-1
0.5B -0.126 -0.017 -0.135
1.5B -0.126 -0.054 -0.188
3B -0.148 -0.158 -0.112
7B -0.114 -0.112 -0.034
Table 6: Correlations between metrics and generation quality (ROUGE-1) on Pref-LaMP. All correlations are negligible to negative across model scales, demonstrating no standard metric predicts behavioral alignment.
Table 6 exposes the complete failure of standard evaluation: RM accuracy, policy accuracy, and win rates show negligible to negative correlations with generation quality (Kendall’s $\tau$ = -0.188 to -0.017). Even policy accuracy—measuring discrimination under the actual deployment mechanism—cannot predict behavioral alignment. Win rates are often most negative, confirming reward hacking.
The takeaway is stark: no existing metric predicts whether methods will generate aligned outputs. Ground-truth evaluation on user-authored completions is not optional—it’s necessary.
9 Discussion and Recommendations
Our findings reveal a fundamental evaluation crisis: RM accuracy (Kendall’s $\tau$ = 0.08-0.31 with policy performance) and discrimination ability both fail to predict generation quality. Methods with 20-point RM accuracy gaps produce identical outputs, while reward models claiming 99% win rates show no ground-truth improvement.
For practitioners: Use ICL-RAG for models $\geq$3B. Reward modeling adds complexity without benefit at scale.
For researchers: (1) Report generation quality on ground-truth completions, not just RM accuracy. (2) Evaluate across model scales to detect scale-dependent failures. (3) Compare against ICL baselines. (4) Build behavioral benchmarks like Pref-LaMP with user-authored completions.
The discrimination-generation gap reflects RGD’s nature as a complex dynamical system. At each generation step, two forces interact: base model logits and reward scores. These can agree or conflict, with dominance shifting across token positions. Policy accuracy measures whether the scoring function on average prefers tokens in chosen responses, but generation is sensitive to outliers—a single poorly-scored token can derail trajectories (causing length changes, topic drift). Moreover, discrimination and generation are distinct objectives: successfully ranking complete sequences doesn’t guarantee steering generation away from undesirable regions of the output space. Understanding these dynamics is critical future work.
Limitations: We focus on inference-time methods (RGD) because it represents the only scalable deployment path—per-user RL fine-tuning remains computationally infeasible for realistic user populations. While RL-based personalization might exhibit different failure modes, it cannot address the fundamental problem we expose: practitioners need to select reward models for deployment, and current metrics fail to predict performance under the only viable deployment mechanism. Generalization beyond our three datasets requires investigation across diverse personalization dimensions (dialogue, values, humor). Our theoretical explanations of the discrimination-generation gap remain preliminary and warrant deeper mechanistic study.
10 Conclusion
We have demonstrated that personalized alignment from preference data faces a fundamental evaluation crisis: standard reward model accuracy metrics do not predict deployment performance under reward-guided decoding. Through systematic evaluation across three datasets and four model scales, we show that (1) RM accuracy correlates only weakly with discrimination ability, (2) discrimination ability doesn’t predict generation quality, and (3) simple in-context learning dominates reward-guided methods for models $\geq$3B parameters.
These findings expose that the field has been optimizing proxy metrics divorced from actual deployment objectives. Current personalized alignment methods fundamentally fail to operationalize user preferences into behavioral adaptation under realistic deployment constraints. The introduction of Pref-LaMP—the first benchmark with ground-truth user completions—enables non-circular evaluation and reveals the extent of this failure through direct behavioral assessment.
For practitioners, the path forward is clear: use ICL with retrieval for personalized generation at scale. For researchers, our work highlights the urgent need for ground-truth behavioral evaluation, theoretical understanding of the discrimination-generation gap, and development of methods that can effectively translate preferences into generation guidance. Only by measuring what actually matters—whether models generate outputs aligned with user preferences—can the field make progress on truly personalized alignment.
Appendix A Extended Related Work and Limitations
Incomparable Accuracy Metrics RLHF and DPO both use pairwise preference accuracy, but these metrics measure fundamentally different things. In RLHF, reward model accuracy measures how well $R_{\theta}$ ranks response pairs. However, the reward model is not the final artifact—it guides a policy through ARGS or RL fine-tuning. The critical question is: does the resulting policy generate aligned responses? Reward model accuracy cannot answer this. A reward model might perfectly rank static pairs while the derived policy fails to generate appropriate responses. In DPO, policy accuracy measures whether $\pi_{\phi}$ assigns higher probability to preferred responses, but only at the likelihood level—not generation quality. These metrics are not comparable across methods, and neither directly measures the ultimate goal: whether generated outputs align with user preferences.
Circular Evaluation Under Frozen Rewards A common RLHF practice adapts policies using reward models, then evaluates by measuring if adapted policies achieve higher rewards than baselines. This creates circularity: the reward model serves as both the training signal and evaluation metric. High performance only confirms the policy learned to exploit the reward model’s scoring function—not that it captures actual user preferences. If the reward model is misspecified, this circular evaluation systematically hides the failure. A policy could achieve high reward scores while generating responses users would disprefer, and the evaluation cannot detect this because both training and evaluation use the same potentially-flawed reward model.
Proxy-Based Evaluation with LLM-as-a-Judge Recent work uses frontier LLMs as judges, conditioning them on few-shot user examples to rank policy outputs. While appealing, LLM-as-judge remains a learned proxy, not a direct measure of user satisfaction. It provides only relative rankings between methods and cannot quantify whether even the best-ranked method produces satisfactory outputs for individual users.
Toward Comprehensive Evaluation These limitations motivate our evaluation framework, which: (1) introduces comparable metrics for both reward model quality and policy understanding, (2) breaks circular evaluation by measuring behavioral alignment against ground-truth user completions rather than reward scores, (3) moves beyond proxies to evaluate actual generation quality, and (4) disentangles where personalization succeeds or fails across the reward modeling, policy guidance, and generation stages.
Appendix B Raw Results: PRISM
Please note that we only evaluate on a subset of the test split of PRISM. This is because policy accuracy computation was expensive. Reward model’s performance on the full test split is in Table 8. Global RM still outperforms all other methods so our conclusions in the main paper text does not change. Meanwhile, data plotted in Figure 1 can be found in Table 7.
LoRE PReF VPL Global RM Global RM-v2 MPU MPU-Avg LoRE-Alt
RM Accuracy 65.95 $\pm$ 4.57 74.34 $\pm$ 2.20 68.35 $\pm$ 2.59 77.94 $\pm$ 3.55 62.11 $\pm$ 4.15 52.76 $\pm$ 5.40 60.43 $\pm$ 1.25 67.15 $\pm$ 0.42
Policy accuracies
0.5B 56.83 $\pm$ 0.00 57.31 $\pm$ 1.66 66.43 $\pm$ 3.24 64.51 $\pm$ 0.83 62.59 $\pm$ 5.04 59.47 $\pm$ 2.91 57.55 $\pm$ 3.81 58.51 $\pm$ 0.83
1.5B 57.07 $\pm$ 0.42 58.99 $\pm$ 1.90 65.71 $\pm$ 1.81 64.51 $\pm$ 0.83 62.35 $\pm$ 4.63 60.43 $\pm$ 2.88 58.27 $\pm$ 2.49 57.55 $\pm$ 0.72
3B 57.07 $\pm$ 0.42 57.31 $\pm$ 0.42 64.03 $\pm$ 3.81 64.51 $\pm$ 0.83 61.39 $\pm$ 3.69 57.55 $\pm$ 2.88 56.12 $\pm$ 2.59 56.59 $\pm$ 1.81
7B 57.55 $\pm$ 0.00 59.23 $\pm$ 2.53 63.79 $\pm$ 2.31 64.51 $\pm$ 0.83 61.87 $\pm$ 3.81 58.03 $\pm$ 2.91 56.35 $\pm$ 2.53 57.07 $\pm$ 0.83
Table 7: Reward model accuracy and policy accuracies per model and method for the PRISM dataset.
LoRE LoRE-Alt MPU MPU-Avg PAL PReF VPL Global RM Global RM-v2
RM Accuracy $56.53$ $\pm 0.15$ $60.33$ $\pm 0.57$ $49.66$ $\pm 1.13$ $49.07$ $\pm 1.46$ $62.51$ $\pm 0.30$ $62.73$ $\pm 0.64$ $60.31$ $\pm 2.98$ $64.51$ $\pm 0.41$ $59.18$ $\pm 0.74$
Table 8: Test set accuracy evaluated across all samples from unseen users. Note: The results reported in the main text use only a single test sample per user due to computational constraints.
Appendix C Raw Results: Pref-LaMP5
Raw results for Pref-LaMP5 dataset can be found in Tables 9, 10, 11 and 12.
Method RM Acc Policy Acc ROUGE-1 ROUGE-L BertScore-F1
Reward Models + ARGS
GenARM 83.89 $\pm$ 0.60 71.62 $\pm$ 0.85 40.82 $\pm$ 0.50 35.13 $\pm$ 0.42 68.93 $\pm$ 0.25
LoRE 65.60 $\pm$ 7.89 67.84 $\pm$ 0.50 41.40 $\pm$ 0.01 35.86 $\pm$ 0.01 68.45 $\pm$ 0.01
LoRE-Alt 84.96 $\pm$ 0.48 64.63 $\pm$ 5.61 35.14 $\pm$ 10.81 30.61 $\pm$ 9.07 67.10 $\pm$ 2.34
MPU 65.26 $\pm$ 1.03 60.96 $\pm$ 8.54 39.97 $\pm$ 2.35 34.43 $\pm$ 2.30 67.69 $\pm$ 1.26
MPU-Avg 67.30 $\pm$ 0.30 60.14 $\pm$ 5.50 41.38 $\pm$ 0.03 35.83 $\pm$ 0.02 68.46 $\pm$ 0.01
PAL 53.77 $\pm$ 0.28 nan $\pm$ nan nan $\pm$ nan nan $\pm$ nan nan $\pm$ nan
PReF 51.46 $\pm$ 3.84 66.06 $\pm$ 2.55 40.88 $\pm$ 0.90 35.26 $\pm$ 0.97 68.14 $\pm$ 0.56
VPL 43.63 $\pm$ 3.77 63.58 $\pm$ 4.02 31.04 $\pm$ 14.71 26.68 $\pm$ 12.78 64.09 $\pm$ 5.90
Vanilla RM 84.96 $\pm$ 0.13 66.41 $\pm$ 0.13 41.37 $\pm$ 0.02 35.87 $\pm$ 0.02 68.45 $\pm$ 0.00
Vanilla RM-v2 84.69 $\pm$ 0.07 79.13 $\pm$ 0.57 31.41 $\pm$ 0.75 26.40 $\pm$ 0.79 67.52 $\pm$ 0.56
DPO
P-DPO - 77.57 $\pm$ 0.92 40.65 $\pm$ 0.46 34.56 $\pm$ 0.43 68.04 $\pm$ 0.24
Baselines
Zero-Shot - - 40.32 $\pm$ 0.00 34.77 $\pm$ 0.00 67.70 $\pm$ 0.00
ICL - nan $\pm$ nan 38.17 $\pm$ 0.20 33.52 $\pm$ 0.21 67.14 $\pm$ 0.13
ICL-RAG - - 39.95 $\pm$ 0.02 35.62 $\pm$ 0.02 68.21 $\pm$ 0.00
Table 9: ROUGE scores for Qwen2.5-0.5B-Instruct (mean $\pm$ std across seeds).
Method RM Acc Policy Acc ROUGE-1 ROUGE-L BertScore-F1
Reward Models + ARGS
GenARM 83.89 $\pm$ 0.60 71.24 $\pm$ 0.52 43.54 $\pm$ 0.74 37.53 $\pm$ 0.79 69.99 $\pm$ 0.04
LoRE 65.60 $\pm$ 7.89 68.49 $\pm$ 0.26 45.47 $\pm$ 0.03 39.45 $\pm$ 0.04 70.32 $\pm$ 0.01
LoRE-Alt 84.96 $\pm$ 0.48 65.11 $\pm$ 6.09 38.55 $\pm$ 11.99 33.50 $\pm$ 10.30 68.69 $\pm$ 2.82
MPU 65.26 $\pm$ 1.03 61.97 $\pm$ 8.37 44.53 $\pm$ 1.59 38.56 $\pm$ 1.46 69.76 $\pm$ 0.90
MPU-Avg 67.30 $\pm$ 0.30 61.15 $\pm$ 6.38 45.47 $\pm$ 0.04 39.45 $\pm$ 0.04 70.31 $\pm$ 0.01
PReF 51.46 $\pm$ 3.84 67.07 $\pm$ 1.74 45.30 $\pm$ 0.39 39.30 $\pm$ 0.35 70.20 $\pm$ 0.18
VPL 43.63 $\pm$ 3.77 64.72 $\pm$ 3.68 35.19 $\pm$ 15.65 30.32 $\pm$ 13.94 66.08 $\pm$ 6.12
Vanilla RM 84.96 $\pm$ 0.13 67.17 $\pm$ 0.04 45.47 $\pm$ 0.02 39.45 $\pm$ 0.03 70.32 $\pm$ 0.00
Vanilla RM-v2 84.69 $\pm$ 0.07 78.73 $\pm$ 0.71 32.75 $\pm$ 0.78 27.77 $\pm$ 0.78 68.27 $\pm$ 0.59
DPO
P-DPO - 77.57 $\pm$ 0.92 40.65 $\pm$ 0.46 34.56 $\pm$ 0.43 68.04 $\pm$ 0.24
Baselines
Zero-Shot - - 43.83 $\pm$ 0.00 37.89 $\pm$ 0.00 69.42 $\pm$ 0.00
ICL - nan $\pm$ nan 43.09 $\pm$ 0.25 37.90 $\pm$ 0.07 69.50 $\pm$ 0.14
ICL-RAG - - 43.96 $\pm$ 0.03 39.16 $\pm$ 0.01 69.96 $\pm$ 0.00
Table 10: ROUGE scores for Qwen2.5-1.5B-Instruct (mean $\pm$ std across seeds).
Method RM Acc Policy Acc ROUGE-1 ROUGE-L BertScore-F1
Reward Models + ARGS
GenARM 83.89 $\pm$ 0.60 71.34 $\pm$ 0.56 43.30 $\pm$ 0.96 37.22 $\pm$ 0.99 69.94 $\pm$ 0.13
LoRE 65.60 $\pm$ 7.89 69.65 $\pm$ 0.31 45.14 $\pm$ 0.03 39.21 $\pm$ 0.02 70.53 $\pm$ 0.01
LoRE-Alt 84.96 $\pm$ 0.48 65.76 $\pm$ 6.60 39.74 $\pm$ 9.42 34.26 $\pm$ 8.65 68.77 $\pm$ 3.07
MPU 65.26 $\pm$ 1.03 63.46 $\pm$ 8.11 44.75 $\pm$ 0.75 38.62 $\pm$ 0.97 70.22 $\pm$ 0.49
MPU-Avg 67.30 $\pm$ 0.30 63.16 $\pm$ 6.70 45.16 $\pm$ 0.05 39.25 $\pm$ 0.03 70.56 $\pm$ 0.04
PReF 51.46 $\pm$ 3.84 68.79 $\pm$ 0.70 44.99 $\pm$ 0.29 39.10 $\pm$ 0.16 70.45 $\pm$ 0.12
VPL 43.63 $\pm$ 3.77 66.75 $\pm$ 2.33 36.00 $\pm$ 14.93 30.90 $\pm$ 13.64 66.71 $\pm$ 6.07
Vanilla RM 84.96 $\pm$ 0.13 68.28 $\pm$ 0.00 45.14 $\pm$ 0.03 39.20 $\pm$ 0.02 70.53 $\pm$ 0.01
Vanilla RM-v2 84.69 $\pm$ 0.07 78.02 $\pm$ 0.94 36.43 $\pm$ 1.99 30.36 $\pm$ 1.77 68.25 $\pm$ 0.43
DPO
P-DPO - 77.57 $\pm$ 0.92 40.65 $\pm$ 0.46 34.56 $\pm$ 0.43 68.04 $\pm$ 0.24
Baselines
Zero-Shot - - 44.86 $\pm$ 0.00 38.67 $\pm$ 0.00 70.28 $\pm$ 0.00
ICL - nan $\pm$ nan 45.63 $\pm$ 0.04 40.14 $\pm$ 0.05 70.47 $\pm$ 0.06
ICL-RAG - - 46.82 $\pm$ 0.03 41.38 $\pm$ 0.04 71.14 $\pm$ 0.00
Table 11: ROUGE scores for Qwen2.5-3B-Instruct (mean $\pm$ std across seeds).
Method RM Acc Policy Acc ROUGE-1 ROUGE-L BertScore-F1
Reward Models + ARGS
GenARM 83.89 $\pm$ 0.60 72.08 $\pm$ 0.58 45.17 $\pm$ 0.53 39.18 $\pm$ 0.56 70.52 $\pm$ 0.11
LoRE 65.60 $\pm$ 7.89 70.97 $\pm$ 0.35 45.97 $\pm$ 0.03 40.16 $\pm$ 0.03 70.75 $\pm$ 0.00
LoRE-Alt 84.96 $\pm$ 0.48 66.52 $\pm$ 7.25 41.07 $\pm$ 8.42 35.66 $\pm$ 7.82 69.10 $\pm$ 2.89
MPU 65.26 $\pm$ 1.03 64.82 $\pm$ 8.31 45.56 $\pm$ 0.66 39.78 $\pm$ 0.66 70.45 $\pm$ 0.47
MPU-Avg 67.30 $\pm$ 0.30 64.02 $\pm$ 6.99 45.92 $\pm$ 0.04 40.14 $\pm$ 0.02 70.75 $\pm$ 0.01
PReF 51.46 $\pm$ 3.84 70.04 $\pm$ 0.45 45.82 $\pm$ 0.18 40.01 $\pm$ 0.23 70.71 $\pm$ 0.08
VPL 43.63 $\pm$ 3.77 67.82 $\pm$ 2.50 37.57 $\pm$ 14.22 32.64 $\pm$ 12.70 67.53 $\pm$ 5.39
Vanilla RM 84.96 $\pm$ 0.13 69.08 $\pm$ 0.04 45.96 $\pm$ 0.04 40.15 $\pm$ 0.02 70.75 $\pm$ 0.01
Vanilla RM-v2 84.69 $\pm$ 0.07 78.44 $\pm$ 0.90 38.91 $\pm$ 4.31 33.43 $\pm$ 3.62 68.96 $\pm$ 0.91
DPO
P-DPO - 77.57 $\pm$ 0.92 40.65 $\pm$ 0.46 34.56 $\pm$ 0.43 68.04 $\pm$ 0.24
Baselines
Zero-Shot - - 45.46 $\pm$ 0.00 39.23 $\pm$ 0.00 70.20 $\pm$ 0.00
ICL - nan $\pm$ nan 47.42 $\pm$ 0.22 42.24 $\pm$ 0.26 71.54 $\pm$ 0.11
ICL-RAG - - 48.90 $\pm$ 0.02 43.81 $\pm$ 0.01 72.20 $\pm$ 0.00
Table 12: ROUGE scores for Qwen2.5-7B-Instruct (mean $\pm$ std across seeds).
Appendix D Architectural and Initialization Enhancements to PReF for Reward-Guided Decoding and Dataset Considerations
D.1 Motivation for Architectural Modification
In our work, we leverage the core principles of PReF but introduce a key architectural modification and a novel initialization scheme. These changes are motivated by the need to adapt PReF from a pairwise preference model into a pointwise reward model, making it suitable for advanced applications such as reward-guided decoding.
The original PReF model is designed to predict a user’s preference for one complete response over another. It computes a single score for a pair of items, $(r_{1},r_{2})$. However, reward-guided decoding requires a scalar reward score for a single, often incomplete, sequence at each step of the generation process. The original PReF formulation is therefore unsuitable for this task.
To address this, we modified the PReF architecture to explicitly compute a user-specific reward for an individual response, $R(u,r)$. This allows us to score single candidate sequences during decoding.
D.2 Original vs. Modified Reward Formulation
The original PReF model calculates a preference score $s$ for a user $u$ and a pair of responses $(r_{1},r_{2})$ based on the difference of their feature representations:
$s(u,r_{1},r_{2})=\mathbf{u}^{\top}\big(\phi(r_{1})-\phi(r_{2})\big)$
Here, $\phi$ is a linear head that projects the LLM’s response embeddings into a latent feature space, and $\mathbf{u}$ is the user’s embedding vector.
Our modified architecture decomposes this calculation into two distinct reward computations:
$R(u,r)=\mathbf{u}^{\top}\phi(r)$
$s(u,r_{1},r_{2})=R(u,r_{1})-R(u,r_{2})$
While these two formulations are mathematically equivalent in the final forward pass, this decomposition presents a significant challenge for the model’s initialization, which we address with a novel technique.
D.3 The Initialization Challenge and Our Solution
The PReF methodology uses Singular Value Decomposition (SVD) on a $(\text{response\_pair}\times\text{user})$ preference matrix to warm-start the model’s parameters. A key challenge arises because the SVD process yields a single feature vector, $\mathbf{v}_{p}$, for each response pair $p$. This vector $\mathbf{v}_{p}$ serves as a proxy for the latent feature difference, $\phi(r_{1})-\phi(r_{2})$.
The SVD provides no direct information about the individual feature vectors $\phi(r_{1})$ and $\phi(r_{2})$. To initialize a linear head $\phi$ that operates on individual responses, we leverage the linearity of the projection and work directly in the difference space.
We achieve this with the following direct regression algorithm:
  1. 1.
    Perform SVD: Perform SVD on the preference matrix as in the original PReF to obtain the matrix of pairwise feature vectors $U_{S}$ (representing $\phi(r_{1})-\phi(r_{2})$ for each pair) and user embeddings $V_{S}$.
  2. 2.
    Compute LLM Embedding Differences: For each unique response pair $(r_{1},r_{2})$ in the training data, compute the difference of their frozen LLM embeddings: $\mathbf{e}_{\text{diff}}=\mathbf{e}(r_{1})-\mathbf{e}(r_{2})$.
  3. 3.
    Direct Regression on Differences: Construct a regression problem where the input matrix $X$ contains all LLM embedding differences $\mathbf{e}_{\text{diff}}$, and the target matrix $Y$ contains the corresponding SVD-derived pairwise features $\mathbf{v}_{p}$ from $U_{S}$. Solve for the linear head weights $W$ such that:
    $W\cdot(\mathbf{e}(r_{1})-\mathbf{e}(r_{2}))\approx\mathbf{v}_{p}$
  4. 4.
    Bias-Free Linear Regression: A bias-free (intercept-free) regression is used to find the optimal initial weights $W$ for the linear head $\phi$. This approach is mathematically sound because:
    • If $\phi(r)=W\cdot\mathbf{e}(r)$, then $\phi(r_{1})-\phi(r_{2})=W\cdot(\mathbf{e}(r_{1})-\mathbf{e}(r_{2}))$
    • Learning $W$ from differences is equivalent to learning it from individual features
    • The lack of bias term means reward values have an arbitrary global offset, which cancels out in pairwise comparisons
D.4 End-to-End Training
Following this warm-start initialization of both the user embeddings (from $V_{S}$) and the linear reward head (from our direct regression method), the model is trained end-to-end using backpropagation. During training, the model computes individual rewards $R(u,r_{1})$ and $R(u,r_{2})$ for chosen and rejected responses. The Bradley-Terry preference learning loss is then computed:
$\mathcal{L}=-\log\sigma(R(u,r_{1})-R(u,r_{2}))$
where $\sigma$ is the sigmoid function. Gradients are backpropagated to fine-tune $\phi$, $\mathbf{u}$, and optionally the LLM encoder if not frozen.
This procedure optimizes the true preference learning objective, with the SVD-based initialization serving as a high-quality starting point that accelerates convergence and improves stability. Any imprecision in initialization (such as the arbitrary offset in absolute reward values) is corrected during training. This enhanced methodology preserves the core insights of PReF’s SVD-based initialization while adapting its architecture to support reward-guided decoding.
D.5 PReF’s Acknowledged Synthetic Augmentation
The original PReF implementation uses a synthetically augmented version of PRISM rather than the natural conversational data. The authors state:
“However, the original PRISM dataset cannot be used directly because it was collected in a way that prevents overlap between users and prompts, which is necessary for our method. Therefore, we augmented it with synthetic annotations via the protocol described in PERSONA, resulting in 50 user preferences per prompt.”
They leverage the PERSONA protocol (TODO cite), which uses LLMs as judges to simulate user preferences. The code snippet confirms this synthetic generation approach:
def get_preference_prism(
                    user_description,
                    prompt,
                    response_1,
                    response_2):
    prompt1=prompts.PRISM_no_confidence.format(
        prompt=prompt,
        user_description=user_description,
        response_1=response_1,
        response_2=response_2,
    pref1 = get_completion(
        prompt1,
        system_prompt=None,
        model="gpt-4o-mini",
        temp=0.0
Code can be found in github codebase idanshen/PReF_code on line 78 in file PReF_code/utils/data.py. This can be found here.
D.6 Data Structure Analysis
Analysis of their dataset reveals a perfectly uniform structure:
Total training samples: 90,450
Unique $(\text{prompt},r_{1},r_{2})$ triples: 1,809
Unique users: 1,200
Users per prompt triple: Exactly 50 (zero variance)
Average samples per user: 75.4
This perfect uniformity demonstrates artificial construction.
D.7 Real PRISM Conversational Data
In contrast, the real PRISM dataset (TODO cite) consists of:
  • Natural multi-turn conversations between users and AI assistants
  • Real human preferences expressed through dialogue
  • Sparse preference matrix (unique conversation contexts)
  • No systematic overlap between users and prompts
Our implementation extracts genuine conversational preferences, yielding $\sim$27K training samples.
D.8 Comparison and Implications
PReF’s “PRISM” dataset differs fundamentally from real PRISM data across multiple dimensions. While PReF uses synthetic data generated by GPT-4o-mini with simulated demographic preferences, real PRISM captures authentic human conversations and actual user choices. The synthetic dataset exhibits a dense matrix structure with exactly 50 users per item, enabling high controllability and strong SVD performance, whereas real PRISM data is characterized by sparse, unique contexts with natural variation that yields weaker SVD results. PReF’s dataset contains over 90K samples compared to 27K in the real data, but this larger volume comes at the cost of realism—the synthetic patterns may not generalize to authentic human behavior in the way that real PRISM’s genuine user interactions do.
D.9 Implications for Reproducibility and Comparison
Key points:
  1. 1.
    Not an apples-to-apples comparison—PReF’s dense setup differs fundamentally from real sparse data.
  2. 2.
    SVD initialization performs better in dense synthetic matrices.
  3. 3.
    Training dynamics differ due to uniform synthetic distribution.
  4. 4.
    Evaluation on simulated preferences may not generalize to real human data.
D.10 Methodological Considerations and Design Choices
PReF’s reliance on dense user-item overlap is intrinsic to collaborative filtering. Sparse real data poses challenges but reflects real-world personalization.
Design options:
  • Match PReF: Use synthetic PERSONA-style preferences for strong SVD initialization.
  • Use Real Data: Accept weaker SVD signals, require stronger regularization and robust training.
  • Hybrid: Augment sparse real data with synthetic overlap.
Our Choice: We prioritize authenticity by using real PRISM conversational preferences in their natural sparse form, tackling the more difficult—but more realistic—personalization problem.

### v2

@@
 Abstract
-Personalized alignment from preference data has focused primarily on improving reward model (RM) accuracy, with the implicit assumption that better preference ranking translates to better personalized behavior.
-However, in deployment, computational constraints necessitate inference-time adaptation via reward-guided decoding (RGD) rather than per-user policy fine-tuning.
-This creates a critical but overlooked requirement: reward models must not only rank preferences accurately but also effectively guide token-level generation decisions.
-We demonstrate that standard RM accuracy fails catastrophically as a selection criterion for deployment-ready personalized alignment.
-Through systematic evaluation across three datasets, we introduce policy accuracy—a metric quantifying whether RGD scoring functions correctly discriminate between preferred and dispreferred responses—and show that RM accuracy correlates only weakly with this discrimination ability (Kendall’s $\tau$ = 0.08–0.31).
-More critically, we introduce Pref-LaMP, the first personalized alignment benchmark with ground-truth user completions, enabling direct behavioral evaluation without circular reward-based metrics.
-On Pref-LaMP, we expose a complete decoupling between discrimination and generation: methods with 20-point RM accuracy differences produce almost identical output quality, and even methods achieving high discrimination fail to generate behaviorally aligned responses.
-Furthermore, simple in-context learning (ICL) dominates all reward-guided methods for models $\geq$3B parameters, achieving 3–5 point ROUGE-1 gains over the best reward method at 7B scale.
-These findings reveal that the field has been optimizing for proxy metrics that don’t predict deployment performance, and that current personalized alignment methods fundamentally fail to operationalize preferences into behavioral adaptation under realistic deployment constraints.
+Personalized alignment from preference data has focused primarily on improving personal reward model (RM) accuracy, with the implicit assumption that better preference ranking translates to better personalized behavior.
+However, in deployment, computational constraints necessitate inference-time adaptation such as reward-guided decoding (RGD) rather than per-user policy fine-tuning.
+This creates a critical but overlooked requirement: reward models must not only rank preferences accurately but also effectively guide generation.
+We demonstrate that standard RM accuracy fails catastrophically as a selection criterion for deployment-ready personalized rewards.
+We introduce policy accuracy—a metric quantifying whether RGD-adapted LLMs correctly discriminate between preferred and dispreferred responses—and show that upstream RM accuracy correlates only weakly with downstream policy accuracy (Kendall’s $\tau$ = 0.08–0.31).
+More critically, we introduce Pref-LaMP, the first personalized alignment benchmark with ground-truth user completions, enabling direct behavioural evaluation.
+On Pref-LaMP, we expose a complete decoupling between discriminative ranking and generation metrics: methods with 20-point RM accuracy differences produce almost identical output quality, and methods with high ranking accuracy can fail to generate behaviorally aligned responses.
+These findings reveal that the field has been optimizing for proxy metrics that do not predict deployment performance, and that current personalized alignment methods fail to operationalize preferences into behavioral adaptation under realistic deployment constraints.
+In contrast, we find simple in-context learning (ICL) to be highly effective - dominating all reward-guided methods for models $\geq$3B parameters, achieving $\sim$3 point ROUGE-1 gains over the best reward method at 7B scale.
 The Reward Model Selection Crisis in Personalized Alignment
-Fady Rezk1,2, Yuangang Pan2, Chuan-Sheng Foo2, Xun Xu2, Nancy Chen2, Henry Gouk1, Timothy Hospedales1,3 1 School of Informatics, University of Edinburgh, UK 2 Agency for Science, Technology and Research (A*STAR), Singapore 3 Samsung AI Research Center, Cambridge, UK
+Fady Rezk1,2, Yuangang Pan2, Chuan-Sheng Foo2, Xun Xu2, Nancy F.
+Chen2, Henry Gouk1, Timothy Hospedales1,3 1 School of Informatics, University of Edinburgh, UK 2 Agency for Science, Technology and Research (A*STAR), Singapore 3 Samsung AI Research Center, Cambridge, UK
 1 Introduction
@@
 This reality motivates personalized alignment: adapting model behavior to heterogeneous, sometimes conflicting, user preferences rather than collapsing them into a single consensus objective.
-Current personalized alignment research has converged on a common paradigm: collect user-specific preference data (pairwise comparisons, ratings, or implicit signals), train personalized reward models to capture individual preferences, and assume that better reward models naturally translates to better policies (Bose et al., 2025; Chen et al., 2025a; Shenfeld et al., 2025; Li et al., 2024; Poddar et al., 2024).
-The last assumption is known to break as reasoned by Goodhart’s law El-Mhamdi and Hoang (2024).
-Unlike standard RLHF, there are no downstream benchmark that measures policy performance such as MMLU Hendrycks et al.
+Current personalized alignment research has converged on a common paradigm: collect user-specific preference data (pairwise comparisons), train personalized ranking/reward models to capture individual preferences, and assume that better reward models naturally translates to better policies (Bose et al., 2025; Chen et al., 2025a; Shenfeld et al., 2025; Li et al., 2024; Poddar et al., 2024).
+The last assumption is likely to break as suggested by Goodhart’s law El-Mhamdi and Hoang (2024).
+Unlike standard RLHF, personalized alignment lacks downstream benchmarks that measures policy performance such as MMLU Hendrycks et al.
 (2021) and GSM8k Cobbe et al.
 (2021).
-1.1 The Deployment Reality and End-to-End Perspective
-Per-user policy fine-tuning using reward model artefacts via RL is computationally infeasible at scale.
-For 1000 user groups, RL-based personalization requires dynamic adapter management, RL instability mitigation, and orders of magnitude more compute than inference-time alternatives.
-The only scalable path is inference-time adaptation through reward-guided decoding (RGD) Khanov et al.
+Practical Deployment and the End-to-End Perspective  Per-user policy fine-tuning using personal rewards via RL is computationally infeasible at scale.
+RL-based personalization requires per-user dynamic adapter management, RL instability mitigation, and orders of magnitude more compute than inference-time alternatives.
+One key scalable deployment path is inference-time adaptation through reward-guided decoding (RGD) Khanov et al.
 (2024), maintaining a single base policy while using personalized rewards to guide generation.
@@
 A personalized alignment method must specify not only how preferences are modeled, but how they are operationalized into behavioral adaptation.
-Papers proposing reward models are responsible for demonstrating that improved RM accuracy translates to improved generation behavior.
-Current evaluation ignores this responsibility, treating reward modeling and policy adaptation as isolated steps.
+A direct corollary of this is that papers proposing reward models are responsible for evaluating if improved RM accuracy translates to improved generation.
+Current evaluations ignore this responsibility, treating reward modeling and policy adaptation as independent.
 This obscures whether methods actually achieve their objective: making models generate responses aligned with user preferences.
-1.2 Our Investigation
-We adopt an end-to-end perspective, studying the complete chain from preference modeling to generation behavior.
-We ask three questions: (1) Does RM accuracy predict discrimination under RGD?
-(2) Does discrimination predict generation quality?
-(3) How do reward methods compare to simpler baselines?
-To answer these, we introduce policy accuracy—measuring whether RGD scoring assigns higher scores to preferred responses—and Pref-LaMP, a benchmark with ground-truth user completions enabling direct behavioral evaluation.
-Our findings reveal a fundamental selection crisis: methods with 20-point RM accuracy differences produce identical generation quality, and correlations between RM accuracy and downstream performance are negligibly weak (Kendall’s $\tau$ = 0.08-0.31).
-Practitioners cannot reliably choose deployment-ready methods because standard metrics don’t predict actual performance.
-1.3 Key Findings
-Finding 1: RM accuracy doesn’t predict discrimination (Kendall’s $\tau$ = 0.08–0.31). Methods with 20-point RM accuracy differences achieve nearly identical policy performance.
-Finding 2: Discrimination doesn’t predict generation. On Pref-LaMP, methods cluster tightly in generation quality despite dramatic variance in RM and policy accuracy.
-Finding 3: ICL dominates at scale. At 3B-7B parameters, ICL-RAG outperforms all reward methods by 3-5 ROUGE-1 points.
-Implications: For practitioners, use ICL-RAG at scale.
-For researchers, RM accuracy is broken—evaluate on ground-truth completions.
-The field needs behavioral benchmarks like Pref-LaMP, analogous to GSM8K/MMLU for general RLHF.
-1.4 Contributions
-We (1) formalize the RM selection problem and introduce policy accuracy for discrimination measurement, (2) contribute Pref-LaMP—the first benchmark with ground-truth user completions, (3) demonstrate RM accuracy fails as a selection criterion across three datasets and four scales, and (4) provide actionable recommendations for practitioners and researchers.
+Our Investigation  We adopt an end-to-end perspective, studying the complete chain from preference modeling to generation behavior.
+We ask three questions: (1) Does RM accuracy predict policy accuracy under RGD?
+(2) Does policy accuracy predict generation quality?
+(3) How do reward-based alignment methods compare to simpler baselines?
+To answer these, we introduce (A) policy accuracy, measuring whether RGD scoring assigns higher scores to preferred responses, and (B) Pref-LaMP, a preference learning benchmark with ground-truth user completions enabling direct behavioral evaluation.
+Key Findings  Our findings reveal a fundamental selection crisis: practitioners cannot reliably choose deployment-ready methods because standard metrics do not predict actual performance.
+Finding 1: Upstream RM accuracy does not predict downstream policy accuracy under RGD (Kendall’s $\tau$ = 0.08–0.31).
+Methods with 20-point RM accuracy differences achieve nearly identical policy performance.
+Finding 2: Response ranking quality does not predict response generation quality. On Pref-LaMP, methods with similar generation quality vary dramatically in RM and policy accuracy.
+Finding 3: ICL dominates at scale. At 7B parameters, ICL-RAG, with RAG selected preference demonstrations Salemi et al. (2024), outperforms best personalized reward model by $\sim$3 ROUGE-1 points.
+Implications: For practitioners, use simple ICL-RAG in preference to published personal reward methods.
+For researchers, take an end-to-end perspective: co-design and co-evaluate reward personalization with policy adaptation strategies and evaluate generation quality as well as ranking.
+Use Pref-LaMP and develop more benchmarks with ground-truth completions, analogous to GSM8K/MMLU for general RLHF.
+Contributions  To summarize, we (1) contribute Pref-LaMP—the first benchmark with ground-truth user completions, (2) demonstrate the standard RM accuracy metric fails as a selection criterion across three datasets and four scales, (3) demonstrate that a simple ICL baseline outperforms published personal alignment work in end-to-end adaptation, and (4) provide actionable recommendations for practitioners and researchers.
 2 Related Work
-2.1 Personalized Alignment
-Recent work has focused on learning user-specific reward models or policies for alignment under limited supervision (e.g., PAL, PReF, LoRE, P-DPO, VPL) (Chen et al., 2025a; Poddar et al., 2024; Bose et al., 2025; Shenfeld et al., 2025; Li et al., 2024).
-These approaches largely target reward-modeling accuracy or relative preference ranking as proxies for personalization quality Chen et al.
+Personalized Alignment  Recent work has focused on learning user-specific reward models or policies for alignment under limited supervision (e.g., PAL, PReF, LoRE, P-DPO, VPL) (Chen et al., 2025a; Poddar et al., 2024; Bose et al., 2025; Shenfeld et al., 2025; Li et al., 2024).
+These approaches largely target reward-modeling accuracy (e.g relative preference ranking) as proxies for personalization quality Chen et al.
 (2025a); Bose et al.
@@
 (2025).
-However, such metrics often fail to capture whether a policy’s generated responses truly reflect an individual user’s preferences, especially under resource-constrained settings or whether it is an unintended overoptimization (Pan et al., 2022).
-2.2 Multi-Objective Alignment
-Multi-objective alignment (MOA) addresses the challenge of optimizing language models across multiple known and predefined reward dimensions simultaneously.
+However, such metrics often fail to capture (i) whether RM adaptation translates to downstream policy adaptation, and (ii) whether, under realistic resource-constrained settings, a personalized policy is able to go beyond response ranking and actually generate responses reflective of a user’s preferences.
+This evaluation limitation leaves open the null hypothesis that prior personal alignment results, measured by RM accuracy, are due to unintended overoptimization, also known as reward-hacking (Pan et al., 2022).
+Multi-Objective Alignment  Multi-objective alignment (MOA) addresses the challenge of optimizing language models across multiple known and predefined reward dimensions simultaneously.
 Unlike personalized alignment, where the goal is to learn individual user preferences under limited supervision, MOA assumes access to distinct reward models for each objective dimension (e.g., helpfulness, harmlessness, factuality) and focuses on finding optimal policy trade-offs among these objectives.
@@
 (2025b) among other methods (Yang et al., 2024b).
-Despite the terminological overlap with ”personalization” in some methods, the key distinction is that MOA operates over known objective dimensions with pre-trained reward models, whereas personalized alignment learns latent user-specific preferences from sparse behavioral data across dataset specific dimensions (e.g., style, length, political values, rhetoric, culture sensitivities, etc).
-2.3 Evaluation Challenges in RLHF
-Existing evaluation practices in RLHF and personalized alignment rely on proxy metrics such as reward-model scores which are susceptible to reward hacking and circularity (Tien et al., 2023; Pan et al., 2022).
-These methods assess optimization success rather than behavioral fidelity Wen et al.
+Evaluation Challenges in RLHF  Existing evaluation practices in RLHF and personalized alignment rely on proxy metrics such as reward-model scores which are susceptible to reward hacking and circularity (Tien et al., 2023; Pan et al., 2022).
+These methods assess optimization success rather than behavioral quality Wen et al.
 (2025); Gao et al.
@@
 Our work addresses these limitations by introducing direct behavioral evaluation on ground-truth user completions, measuring actual generation quality rather than relying on proxy metrics.
-Extended discussion of evaluation pathologies appears in Appendix A.
-2.4 Inference-Time Alignment
-Reward-guided decoding (ARGS) Khanov et al.
+Inference-Time Alignment  Reward-guided decoding Khanov et al.
 (2024) and Best-of-N sampling Ichihara et al.
 (2025) enable policy steering without fine-tuning, making them computationally attractive for personalization.
-Recent work has explored their effectiveness Wu (2025), but evaluation remains limited to reward-based metrics.
-Our work is the first to systematically evaluate these methods for personalized alignment with ground-truth behavioral assessment, revealing fundamental limitations in their ability to operationalize user preferences.
+Recent work has explored their effectiveness Wu (2025), but standard personal alignment evaluation remains limited to reward-based metrics.
+Our work is the first to systematically evaluate test-time alignment (reward-guided decoding in particular) for personalized alignment with ground-truth behavioral assessment, revealing fundamental limitations in their ability to operationalize user preferences.
 3 Preliminaries and Problem Setup
 3.1 Personalized Preference Learning
-Consider a preference dataset $\mathcal{D}=\{(u_{i},x_{i},y_{i}^{(c)},y_{i}^{(r)})\}_{i=1}^{N}$, where $u_{i}\in U=\{1..K\}$ is a user identifier, $x_{i}$ is a prompt, and $y_{i}^{(c)},y_{i}^{(r)}$ are chosen and rejected completions.
-We split users into training $\mathcal{U}_{\mathrm{train}}$ (learning shared structure) and adaptation $\mathcal{U}_{\mathrm{adapt}}$ (few-shot personalization at deployment).
-For user $k$, define $\mathcal{D}_{k}=\{(x_{i},y_{i}^{(c)},y_{i}^{(r)}):u_{i}=k\}$.
+Consider a preference dataset $\mathcal{D}=\{(u_{i},x_{i},y_{i}^{(w)},y_{i}^{(l)})\}_{i=1}^{N}$, where $u_{i}\in\{1..K\}$ is a user identifier, $x_{i}$ is a prompt, and $y_{i}^{(w)},y_{i}^{(l)}$ are chosen/winning and rejected/loosing completions.
+We partition users into $\mathcal{U}_{\mathrm{train}}$ (for learning shared preference structure) and $\mathcal{U}_{\mathrm{adapt}}$ (for evaluating few-shot personalization).
+Users in $\mathcal{U}_{\mathrm{adapt}}$ are further split into support sets $\mathcal{D}_{k}^{\mathrm{support}}$ (for adaptation) and query sets $\mathcal{D}_{k}^{\mathrm{query}}$ (for evaluation).
+For user $k$, we denote their full dataset as $\mathcal{D}_{k}=\{(x_{i},y_{i}^{(w)},y_{i}^{(l)}):u_{i}=k\}$.
 Personalized Reward Modeling.
@@
 Training on $\mathcal{U}_{\mathrm{train}}$ learns both $\theta$ and $\{z_{k}\}_{k\in\mathcal{U}_{\mathrm{train}}}$.
-At deployment, for new user $k\in\mathcal{U}_{\mathrm{adapt}}$ with few-shot data $\mathcal{D}_{k}^{\mathrm{few-shot}}$, we adapt:
-$z_{k}=\mathcal{A}_{R}\big(\mathcal{D}_{k}^{\mathrm{few-shot}};\theta\big)$
-where $\mathcal{A}_{R}$ is the adaptation algorithm.
+At deployment, for user $k\in\mathcal{U}_{\mathrm{adapt}}$ with support data $\mathcal{D}_{k}^{\mathrm{support}}$, we adapt:
+$z_{k}=\mathcal{A}\big(\mathcal{D}_{k}^{\mathrm{support}};\theta\big)$
+where $\mathcal{A}$ is the adaptation algorithm.
 3.2 Deployment via Reward-Guided Decoding
-Given computational constraints prohibiting per-user policy fine-tuning, we deploy personalized alignment through inference-time guidance. We consider two methods:
-Reward-Guided Decoding (ARGS). At each generation step $t$, ARGS scores the top-$k$ tokens $V^{(k)}_{t}$ as:
+Given computational constraints prohibiting per-user policy fine-tuning, we deploy personalized alignment through inference-time guidance using Reward-Guided Decoding Khanov et al. (2024).
+Reward-Guided Decoding (ARGS).  At each generation step $t$, ARGS scores the top-$k$ tokens $V^{(k)}_{t}$ retrieved from an off-the-shelf prior LLM policy, $\pi$, using
 $\text{score}(v\mid x,y_{<t};z_{k},\lambda)$ $=\log\pi(v\mid x,y_{<t})$ (1)
-$\quad+\lambda\cdot r_{\theta,z_{k}}(v\mid x,y_{<t})$
+$\quad+\lambda\cdot r_{\theta,z_{k}}(v\mid x,y_{<t}),$
 selecting $y_{t}=\operatorname*{arg\,max}_{v\in V^{(k)}_{t}}\text{score}(v)$. This balances base model fluency ($\log\pi$) with personalized alignment ($r_{\theta,z_{k}}$).
 3.3 The Evaluation Gap
-Standard practice evaluates personalization methods by reward model accuracy:
+Standard practice evaluates personalization methods by reward model accuracy, defined below.
 Definition 1 (Reward Model Ranking Accuracy).
-For user $k$ with evaluation set $\mathcal{D}_{k}^{\text{eval}}=\{(x_{i},y_{i}^{(c)},y_{i}^{(r)})\}$:
-$\frac{1}{|\mathcal{D}_{k}^{\text{eval}}|}\sum_{\mathcal{D}_{k}^{\text{eval}}}\mathbb{I}\Big[r_{\theta,z_{k}}(y_{i}^{(c)}\mid x_{i})>r_{\theta,z_{k}}(y_{i}^{(r)}\mid x_{i})\Big]$ (2)
-This measures pairwise ranking on complete responses.
-However, deployment via RGD requires token-level scoring.
-A reward model achieving high sequence-level ranking accuracy may provide poor token-level guidance, yet standard evaluation cannot detect this mismatch.
-4 Policy Accuracy: Measuring Discrimination Under RGD
-To bridge the evaluation gap, we introduce a metric quantifying whether the RGD scoring function—not just the reward model in isolation—can discriminate preferred from dispreferred responses.
+For user $k$ with evaluation set $\mathcal{D}_{k}^{\text{query}}=\{(x_{i},y_{i}^{(w)},y_{i}^{(l)})\}$, we define the Reward Model Ranking Accuracy as
+$\frac{1}{|\mathcal{D}_{k}^{\text{eval}}|}\sum_{\mathcal{D}_{k}^{\text{eval}}}\mathbb{I}\Big[r_{\theta,z_{k}}(y_{i}^{(w)}\mid x_{i})>r_{\theta,z_{k}}(y_{i}^{(l)}\mid x_{i})\Big].$ (2)
+This measures pairwise ranking on complete responses of the reward model. Later we will show that this standard metric has several issues and is not predictive of deployment performance.
+4 Policy Accuracy: Measuring Preference Ranking Under RGD
+We introduce a metric quantifying whether the RGD scoring function—not just the reward model in isolation—correctly ranks preferred over dispreferred responses.
 Definition 2 (Policy Ranking Accuracy).
-Let $s_{\pi}:\mathcal{Y}\times\mathcal{X}\to\mathbb{R}$ be the policy scoring function used at generation time. The policy accuracy for user $k$ is:
-$\frac{1}{|\mathcal{D}_{k}^{\text{eval}}|}\sum_{\mathcal{D}_{k}^{\text{eval}}}\mathbb{I}\Big[s_{\pi}(y_{i}^{(c)}\mid x_{i})>s_{\pi}(y_{i}^{(r)}\mid x_{i})\Big]$ (3)
-We instantiate $s_{\pi}$ with three scoring functions, each revealing different aspects of the personalization pipeline:
-Prior Policy. The base policy’s length-normalized log-likelihood, quantifying zero-shot discrimination ability:
-$\frac{1}{|y|}\sum_{t=1}^{|y|}\log\pi(y_{t}\mid x,y_{<t})$ (4)
-Global RM Posterior. RGD scoring with a non-personalized reward model:
-$\sum_{t=1}^{|y|}\Big[\log\pi(y_{t}\mid x,y_{<t})+\lambda\cdot r_{\theta}(y_{t}\mid x,y_{<t})\Big]$ (5)
-Personalized Posterior. RGD scoring with personalized reward model:
-$\sum_{t=1}^{|y|}\Big[\log\pi(y_{t}\mid x,y_{<t})$ $\quad+\lambda\cdot r_{\theta,z_{k}}(y_{t}\mid x,y_{<t})\Big]$ (6)
-Comparing these accuracies reveals: (1) whether reward guidance improves discrimination over the base policy, and (2) whether personalization provides gains beyond global reward models.
-Critical Limitation.
-Policy accuracy measures discrimination ability—whether the scoring function prefers chosen responses—not generation quality.
-High policy accuracy provides no guarantee the policy will generate outputs that behaviorally align with user preferences.
-A method might correctly rank static responses while producing generations that differ substantially from what users would actually write.
-This motivates our ground-truth behavioral evaluation in Section 5.
+Let $s:\mathcal{Y}\times\mathcal{X}\to\mathbb{R}$ be the scoring function used at generation time. The policy accuracy for user $k$ is given by
+$\frac{1}{|\mathcal{D}_{k}^{\text{query}}|}\sum_{\mathcal{D}_{k}^{\text{query}}}\mathbb{I}\Big[s(y_{i}^{(w)}\mid x_{i})>s(y_{i}^{(l)}\mid x_{i})\Big],$ (3)
+where $y^{(w)}$ and $y^{(l)}$ denote the chosen (winning) and rejected (losing) completions.
+We instantiate $s$ with three scoring functions, each revealing different aspects of the personalization pipeline.
+Base Policy. The base policy’s length-normalized log-likelihood, its off-the-shelf non-personalized zero-shot ranking ability,
+$s_{\text{base}}(y\mid x)=\frac{1}{|y|}\sum_{t=1}^{|y|}\log\pi(y_{t}\mid x,y_{<t}).$ (4)
+Global RGD. RGD with a non-personalized reward model $r_{\theta}$ trained by aggregating data across all users denoted $s_{\text{global}}(y\mid x)$,
+$\sum_{t=1}^{|y|}\Big[\log\pi(y_{t}\mid x,y_{<t})+\lambda\cdot r_{\theta}(y_{t}\mid x,y_{<t})\Big].$ (5)
+Personalized RGD. RGD with personalized reward model $r_{\theta,z_{k}}$ denoted $s_{\text{personal}}(y\mid x;z_{k})$,
+$\sum_{t=1}^{|y|}\Big[\log\pi(y_{t}\mid x,y_{<t})+\lambda\cdot r_{\theta,z_{k}}(y_{t}\mid x,y_{<t})\Big].$ (6)
+Comparing these reveals: (1) whether reward guidance improves ranking over the base policy, and (2) whether personalization provides gains beyond a global reward model.
+In-Context Learning.
+An alternative personalization mechanism conditions the base policy on user-specific demonstrations $\mathcal{D}_{k}^{\text{demo}}=\{(x_{i},y_{i}^{(w)},y_{i}^{(l)}\}\subset\mathcal{D}_{k}^{\text{support}}$ rather than learning reward parameters.
+The ICL scoring function, denoted $s_{\text{ICL}}(y\mid x;\mathcal{D}_{k}^{\text{demo}})$ is
+$\frac{1}{|y|}\sum_{t=1}^{|y|}\log\pi(y_{t}\mid\mathcal{D}_{k}^{\text{demo}},x,y_{<t}),$ (7)
+[…]
+Outstanding Limitation.
+Policy accuracy measures how well the scoring function ranks static responses—not whether the policy will actually generate outputs that actually align with user preferences.
+A method might rank existing responses correctly while producing generations that differ substantially from what users would write.
+This motivates our behavioral evaluation in Section 5.
 5 Pref-LaMP: A Benchmark for Direct Behavioral Evaluation
 To enable direct measurement of behavioral alignment without circular reward-based metrics, we introduce Pref-LaMP—a personalized alignment benchmark providing both pairwise preferences and ground-truth user-authored completions.
-5.1 Dataset Construction
-Pref-LaMP derives from LaMP-5 Salemi et al.
+Dataset ConstructionPref-LaMP derives from LaMP-5 Salemi et al.
 (2024), pairing researchers’ abstracts with their titles.
 Both are author-written, capturing individual style.
-We construct preferences via hard negative mining: (1) encode abstracts with Qwen3-Embedding-0.6B, (2) retrieve top-$k$ similar abstracts from other users, (3) sample one retrieved abstract and use its title as $y^{(r)}$, (4) use original title as $y^{(c)}$.
-This ensures rejections are topically relevant but stylistically different.
-Human-written negatives avoid shortcut learning—initial LLM-generated rejections let linear probes detect generation artifacts rather than preference signals.
-Pref-LaMP is the first benchmark enabling non-circular evaluation through user-authored completions, measurable via ROUGE and BERTScore.
-5.2 Behavioral Alignment Metric
+We construct preferences via hard negative mining: (1) encode abstracts with Qwen3-Embedding-0.6B, (2) retrieve top-$k$ similar abstracts, (3) sample one retrieved abstract as $x$ and use its title as $y^{(l)}$, (4) use original title as $y^{(w)}$.
+This ensures rejections are topically relevant but different in title formulation [^1] Human-written rather than LLM-written negatives avoid shortcut learning.
+Initial LLM-generated rejections let linear probes detect generation artifacts rather than preference signals..
+Pref-LaMP is the first benchmark enabling direct behavioural evaluation of personalization through user-authored completions, measurable via ROUGE and BERTScore.
+Behavioral Alignment Metric  We evaluate end-to-end behavioural alignment by comparing user-generated and personalized model responses.
 Definition 3 (Behavioral Alignment).
-Let $\mathcal{G}:\mathcal{X}\to\mathcal{Y}$ be a generation operator and $\mathcal{S}:\mathcal{Y}\times\mathcal{Y}\to[0,1]$ be a similarity measure.
+Let $\mathcal{G}:\mathcal{X}\to\mathcal{Y}$ be a generation operator and $\mathcal{S}:\mathcal{Y}\times\mathcal{Y}\to\mathcal{R}$ be a similarity measure.
 For user $k$ with test set $\mathcal{P}_{k}=\{(x_{j},y_{j}^{\text{GT}})\}_{j=1}^{M_{k}}$:
@@
 We instantiate $\mathcal{G}$ with ARGS decoding (Eq.
-1), and $\mathcal{S}$ with ROUGE-1 (lexical overlap), ROUGE-L (longest common subsequence), and BERTScore-F1 (semantic similarity).
+1), zero-shot generation and ICL generation.
+Meanwhile, $\mathcal{S}$ in instantiated with ROUGE-1 (lexical overlap), ROUGE-L (longest common subsequence), and BERTScore-F1 (semantic similarity).
 This measures whether generated outputs match what users actually write, providing ground-truth assessment of behavioral alignment.
 6 Experimental Setup
-6.1 Datasets and Models
-TLDR Stiennon et al.
+Datasets and Models  We consider datasets: TLDR Stiennon et al.
 (2020): Binary stylistic preferences, 10 training users (2,097 prefs/user), 31 adaptation users (100 prefs/user).
@@
 Pref-LaMP (ours): User-authored completions, 485 training users (48.8 prefs/user), 126 adaptation users (49.2 prefs/user).
-6.2 Models and Methods
 Dataset Reward Model Base Policies
@@
 Table 1: Model configurations.
-All reward models use LoRA rank 8, trained on $\mathcal{U}_{\text{train}}$ to learn shared $\theta$ and user-specific $\{z_{k}\}$.
+Models and Methods  All reward models use LoRA rank 8, trained on $\mathcal{U}_{\text{train}}$ to learn shared $\theta$ and user-specific $\{z_{k}\}$.
 We evaluate six personalization methods: LoRE Bose et al.
@@
 (2025) (autoregressive RM for token-level guidance), zero-shot generation, ICL (random demonstrations), and ICL-RAG (retrieved demonstrations).
-6.3 Evaluation Protocol
-We measure: (1) RM accuracy on adaptation/deployment users held-out preferences, (2) policy accuracy using prior/global/personalized scoring functions, (3) generation quality on Pref-LaMP via ROUGE-1/L and BERTScore against ground-truth, and (4) win rates where each method’s RM judges its own outputs versus zero-shot baseline.
-7 Results Part 1: Discrimination Does Not Follow From Ranking
-We first investigate whether reward model accuracy predicts discrimination ability under reward-guided decoding—i.e., whether methods that rank preferences accurately can also guide generation appropriately.
-7.1 TLDR: Weak Correlation on Simple Data
+Evaluation Protocol  We measure: (1) RM accuracy on adaptation users’ held-out preferences (Eq.
+2), (2) Adaption users’ policy accuracy vs prior (no-reward) and global reward baselines (Eq.
+3), (3) generation quality on Pref-LaMP via ROUGE-1/L and BERTScore against ground-truth (Eq.
+8), and (4) win rates where each method’s RM judges its own outputs versus zero-shot baseline.
+7 Result 1: Reward Model Accuracy Does Not Imply Policy Accuracy
+We first investigate whether reward model accuracy predicts policy accuracy under reward-guided decoding. I.e., whether personal rewards that rank preferences well can guide policies to do the same.
 Policy Accuracy
 Method RM Accuracy 135M 360M 1.7B
-LoRE 82.56 $\pm$ 1.12 81.09 $\pm$ 0.76 81.09 $\pm$ 0.89 81.01 $\pm$ 1.10
+Personalized LoRE 82.56 $\pm$ 1.12 81.09 $\pm$ 0.76 81.09 $\pm$ 0.89 81.01 $\pm$ 1.10
 MPU 76.76 $\pm$ 0.26 49.96 $\pm$ 3.24 54.36 $\pm$ 7.36 55.12 $\pm$ 8.07
@@
 VPL 68.67 $\pm$ 1.94 59.78 $\pm$ 4.34 60.40 $\pm$ 4.82 60.67 $\pm$ 4.92
+Baselines GenARM 53.67 $\pm$ 3.15 53.06 $\pm$ 0.02 61.56 $\pm$ 0.03 62.59 $\pm$ 0.02
 Global 73.67 $\pm$ 2.45 58.80 $\pm$ 4.79 59.58 $\pm$ 1.55 60.34 $\pm$ 1.80
 Global-V2 38.66 $\pm$ 1.40 53.05 $\pm$ 0.02 61.26 $\pm$ 0.07 62.54 $\pm$ 0.04
-GenARM 53.67 $\pm$ 3.15 53.06 $\pm$ 0.02 61.56 $\pm$ 0.03 62.59 $\pm$ 0.02
 Prior - 62.00 $\pm$ 0.00 61.57 $\pm$ 0.00 64.60 $\pm$ 0.00
@@
 Kendall $\tau$ - 0.31 0.08 0.09
-Table 2: Policy accuracy and reward model performance on TLDR + SmolLM. Base reward model is SmolLM2-135M
-On TLDR’s binary preferences, most personalization methods (MPU, MPU-Avg, PReF) achieve policy accuracies below the prior baseline (62-65%), particularly at smaller scales.
-RM accuracy correlates weakly with policy accuracy (Kendall’s $\tau$ = 0.08–0.31), degrading as scale increases (Pearson $r$: 0.41 $\to$ 0.11 from 180M to 1.7B).
+Table 2: Policy accuracy and reward model performance on TLDR + SmolLM. Base reward model is SmolLM2-135M. Top: Personal rewards. Mid: Global alignment.
+TLDR: Weak Correlation on Simple Data  We first evaluate the popular TDLR dataset’s simple binary style preferences in Table 2.
+The main observation is that upstream RM accuracy correlates weakly with downstream policy accuracy (Kendall’s $\tau$ = 0.08–0.31), degrading as scale increases (Pearson $r$: 0.41 $\to$ 0.11 from 180M to 1.7B).
+Additionally, most personalization methods (MPU, MPU-Avg, PReF) fail to adapt policies, achieving policy accuracies below the prior baseline (62-65%), particularly at smaller scales.
 PAL achieves highest RM accuracy (87.0%) but doesn’t support RGD, while LoRE—with lower RM accuracy (82.6%)—achieves superior policy accuracy ($\sim$81%).
-LoRE substantially outperforms all baselines with 16-20 point gains over prior policy, demonstrating genuine effectiveness and validating our evaluation framework.
-7.2 PRISM: Personalization Fails on Pluralistic Data
+Overall, LoRE substantially outperforms all competitors as well as prior and global reward policies, demonstrating genuine effectiveness for inference-time adaptation and validating our evaluation framework.
+Personalized Global
 LoRE LoRE-Alt MPU MPU-Avg PAL PReF VPL Global RM Global RM-v2
-RM Accuracy $65.95$ $\pm 4.57$ $67.15$ $\pm 0.42$ $52.76$ $\pm 5.40$ $60.43$ $\pm 1.25$ $70.74$ $\pm 0.42$ $74.34$ $\pm 2.20$ $68.35$ $\pm 2.59$ 77.94 $\pm 3.55$ $62.11$ $\pm 4.15$
-Table 3: Reward model accuracies of various algorithms using Qwen2.5-0.5B backbone on PRISM dataset. Global reward model V1 outperforms all other methods.
+RM Acc. $65.95^{\pm 4.57}$ $67.15^{\pm 0.42}$ $52.76^{\pm 5.40}$ $60.43^{\pm 1.25}$ $70.74^{\pm 0.42}$ $74.34^{\pm 2.20}$ $68.35^{\pm 2.59}$ $\textbf{77.94}^{\pm 3.55}$ $62.11^{\pm 4.15}$
+Table 3: Reward model accuracies of various personal alignment methods using Qwen2.5-0.5B backbone on PRISM dataset. All personal alignment methods underperform non-personal Global Reward Model V1.
 Figure 1: RM vs Policy Accuracy Correlation on PRISM across model scales.
@@
 While correlations slightly strengthen with scale, they remain far below what would be needed for RM accuracy to reliably predict policy performance.
-Notably, methods cluster vertically (similar RM accuracy, varying policy performance), demonstrating that high RM accuracy does not guarantee better policy understanding.
-Results are summarized in Table 3 and Figure 1.
-With reference to RM accuracies in Table 3, all personalized methods underperform Global RM (77.9% vs.
-52.8-74.3%), with even the best (PReF, 74.3%) falling 3.6 points short.
+Notably, methods with similar RM accuracy can have substantially different policy accuracy and vice-versa, demonstrating that the standard RM metric is not a reliable proxy for deployment performance.
+PRISM: Personalization Fails on Pluralistic Data  Results for the more complex PRISM benchmark are summarized in Table 3 and Figure 1.
+All personalization methods fail in terms of their policy accuracy underperforming the global RM baseline (77.9% vs.
+52.8-74.3%) in Table 3.
 Meanwhile, RM-policy correlation, as shown in Fig 1, remains weak (Kendall’s $\tau$ = 0.17-0.29), slightly strengthening with scale ($r$: 0.31 $\to$ 0.48).
 Critically, VPL achieves highest policy accuracy (63.8-66.4%) despite 9.5 points lower RM accuracy than Global RM—a complete ranking inversion.
-Methods show minimal scaling gains, remaining in narrow bands (LoRE: 56.8-57.6%, VPL: 63.8-66.4%), suggesting limitations in learning from sparse pluralistic.
+In terms of scale, methods show minimal scaling gains, remaining in narrow bands (LoRE-Alt: 57.1-58.5%, VPL: 63.8-66.4%).
 Unlike TLDR/SmolLM2 where correlations degraded with scale ($r$ = 0.41→0.11), PRISM/Qwen2.5 shows strengthening correlations ($r$ = 0.31→0.48).
 Whether this reflects dataset differences, model architecture, or their interaction remains unclear.
-Regardless, even at 7B, correlations remain too weak for reliable prediction.
-Reconciling with Prior Claims.
-LoRE’s released code fails to reproduce claimed gains over Global-RM [^1] https://github.com/facebookresearch/LoRe/issues/1; our reimplementation of their linear probe setup yields almost identical accuracy (us: 55.59% vs.
-theirs: 55.55%) validating our imlpementation.
-PReF was evaluated on PERSONA, not PRISM which they admittedly fail on (Appendix D).
-PAL and VPL used different datasets (TLDR, non-overlapping tasks).
-Implication: RM accuracy cannot guide deployment—methods with 10+ point gaps perform identically or invert in ranking.
-This implies that RM accuracy does not measure behaviorally-relevant properties for policy guidance.
-Reward models can rank preferences accurately while failing to provide useful token-level guidance.
+Regardless, even at 7B, correlations remain too weak for choosing methods based on RM accuracy.
+Discussion  Our careful control evaluation shows wide failure of prior personal alignment methods both in terms of beating global alignment baselines, and in terms of the standard metric of RM accuracy not corresponding to downstream policy accuracy.
+We attribute this to a mixture of released code not reproducing results, missing non-personal baselines, and inconsistent non-comparable choice of datasets in prior evaluations.
+See Appendix D and E for further discussion.
+Implication: Personal RM accuracy does not reflect performance during policy inference and cannot guide choice of reward model for deployment: Methods with 10+ point RM gaps can perform identically as adapted policies; methods with near identical RM accuracy can have 10+ point gaps in policy accuracy; and alignment methods can invert in ranking between reward and policy evaluations.
 Recommendation.
-Evaluations must assess downstream policy understanding across multiple datasets and scales—not just RM accuracy.
+Future personal alignment methods must specify a policy adaptation strategy, and assess downstream policy understanding across multiple datasets and scales—not just upstream personal reward accuracy.
 The RM-policy disconnect demands new metrics measuring reward models’ suitability for guiding generation, rather than pairwise ranking accuracy alone.
-8 Results Part 2: Discrimination Does Not Imply Generation
-Given the weak correlation between RM accuracy and policy discrimination ability under RGD, we now ask: even when methods achieve high policy accuracy—demonstrably preferring chosen over rejected responses—do they generate outputs that behaviorally align with user preferences?
-8.1 Pref-LaMP: Complete Decoupling Revealed
-LoRE LoRE-Alt MPU MPU-Avg PAL PReF VPL Vanilla Vanilla-v2 GenARM
-RM Acc $65.60$ $\pm 7.89$ $84.96$ $\pm 0.48$ $65.26$ $\pm 1.03$ $67.30$ $\pm 0.30$ $53.77$ $\pm 0.28$ $51.46$ $\pm 3.84$ $43.63$ $\pm 3.77$ $84.96$ $\pm 0.13$ $84.69$ $\pm 0.07$ $83.89$ $\pm 0.60$
-Table 4: Reward model accuracies using Qwen2.5-0.5B on Pref-LaMP. Vanilla RMs again outperform personalization methods.
+8 Result 2: Preference Discrimination Does Not Imply Generation Quality
+Given the weak correlation between RM accuracy and policy discrimination ability under RGD, we now ask: even when methods achieve high policy accuracy—demonstrably preferring chosen over rejected responses—do they actually generate outputs that behaviorally align with user preferences?
+8.1 Pref-LaMP: Complete Decoupling
+Global Personalized
+Acc GenARM Global RM Global RM-v2 LoRE LoRE-Alt MPU MPU-Avg PAL PReF VPL ICL
+RM $83.89^{\pm 0.60}$ $84.96^{\pm 0.13}$ $84.69^{\pm 0.07}$ $65.60^{\pm 7.89}$ $84.96^{\pm 0.48}$ $65.26^{\pm 1.03}$ $67.30^{\pm 0.30}$ $53.77^{\pm 0.28}$ $51.46^{\pm 3.84}$ $43.63^{\pm 3.77}$ -
+Policy $71.34^{\pm 0.56}$ $68.28^{\pm 0.00}$ $78.02^{\pm 0.94}$ $69.65^{\pm 0.31}$ $65.76^{\pm 6.60}$ $63.46^{\pm 8.11}$ $63.16^{\pm 6.70}$ - $8.79^{\pm 0.70} 66$$6.75^{\pm 2.33}$ $75.67^{\pm 0.44}$
+Table 4: Pref-LaMP preference ranking accuracy for RMs and adapted policies (Qwen2.5-3B).
+RM personalization does not clearly outperform global RMs for either upstream RM or downstream policy preference ranking.
 Figure 2: Generation quality (ROUGE-1, ROUGE-L, BERTScore-F1) under RGD across model scales.
 At 0.5B-1.5B, personalized RMs marginally improve over zero-shot; at 3B-7B, ICL baselines dominate all reward-guided methods.
-Global RMs dominate (84.7-85.0%), with GenARM (83.9%) and LoRE-Alt close. Most personalization methods fail (VPL: 43.6%, PReF: 51.5%). Figure 2 reveals three critical failures:
-(1) Architecture provides no advantage. GenARM, designed for token-level guidance, performs comparably to vanilla methods despite similar RM accuracy.
-(2) RM accuracy decouples from generation.
-Best reward methods (LoRE, Vanilla RM, MPU-Avg) achieve identical ROUGE-1 (45.9-46.0) and BERTScore (70.5-70.8) at 7B despite 20-point RM accuracy gaps.
-High RM accuracy guarantees nothing about generation quality.
-(3) Reward guidance fails at scale.
-At 0.5-1.5B, reward methods improve over zero-shot.
-At 3-7B, ICL-RAG dominates, achieving 48.9 ROUGE-1 and 72.2 BERTScore at 7B—3.0 and 1.5 point gaps over best reward methods.
-RGD becomes counterproductive as base capability increases.
+We first study Pref-LaMP with preference ranking evaluation in Table 4.
+They key observation is that for this challenging task, similarly to PRISM (Table 3), personal alignment methods struggle to surpass Global RM baselines – for both the standard proxy metric of upstream RM accuracy, as well as our downstream policy accuracy.
+Only LoRE-Alt come close to the global baselines in RM accuracy.
+We next move to analysing behavioural generation quality of the policies – as uniquely enabled by our Pref-LAMP dataset, in Figure 2 and raw results in Appendix C.
+We see that: (1) The top personal alignment methods for upstream ranking accuracy (LoRE-Alt in Table 4) tend to underperform in downstream generation quality.
+(2) A few personal alignment methods can surpass the zero-shot baseline but produce comparative performance to Global RM baseline.
+However, the better methods for generation (e.g, MPU-Avg and LoRe) are worse for upstream ranking (Table 4).
+Both these observations reflect decoupling between upstream RM accuracy and downstream behavioural generation.
+This shows that downstream generation quality evaluation is a crucial missing component of standard evaluation practice.
+Our end-to-end behavioural evaluation also allows direct comparison between existing RM-focused personal alignment approaches and ICL.
+From Table 4, we can see that ICL actually achieves better policy preference ranking than the personal RMs.
+In terms of generation quality, Figure 2 shows that direct application of ICL surpasses both the baselines and prior personal alignment methods at 3-7B scale.
+This suggests that practitioners today should use simple ICL in favour of complex RM-based alignment approaches.
 Method 0.5B 1.5B 3B 7B
-GenARM 100.0% 100.0% 100.0% 99.8%
-LoRE 73.7% 72.9% 73.8% 73.7%
+Global GenARM 100.0% 100.0% 100.0% 99.8%
+Global RM 99.9% 99.9% 99.9% 99.9%
+Global RM-v2 98.4% 97.2% 89.4% 84.4%
+Personalized LoRE 73.7% 72.9% 73.8% 73.7%
 LoRE-Alt 97.0% 96.0% 96.3% 96.3%
@@
 VPL 51.1% 50.1% 50.5% 48.4%
-Vanilla RM 99.9% 99.9% 99.9% 99.9%
-Vanilla RM-v2 98.4% 97.2% 89.4% 84.4%
 Table 5: Win rate: fraction of examples where each RM judges its own RGD output as better than zero-shot.
-High win rates reveal severe reward hacking—methods claim near-perfect improvement despite ground-truth metrics showing minimal or negative gains.
-Win rates expose reward hacking: GenARM claims 100% improvement over zero-shot despite ROUGE-1 being marginally worse.
-Vanilla RM claims 99.9% superiority while producing outputs identical to methods 20 points lower in RM accuracy.
-Circular evaluation actively misleads—methods optimize their scoring functions then declare success while actual alignment remains unmeasured.
-Figure 3: ROUGE-1 vs. number of in-context examples. ICL-RAG consistently outperforms random ICL, with performance scaling with both model size and context length.
-ICL-RAG improves steadily with demonstrations and scale, reaching $\sim$49 ROUGE-1 at 7B with 5 shots.
-Larger models show no saturation, effectively leveraging context without reward guidance.
-This reveals RGD’s failure is not inherent to personalization but to its operationalization mechanism.
-8.2 No Metric Predicts Generation Quality
+High win rates reveal severe reward hacking—methods claim near-perfect improvement despite ground-truth metrics showing minimal or negative gains (Fig 2).
+Discussion: Reconciling Prior Claims  How can we reconcile prior papers’ claims of successful RM+RGD based non-personal alignment with the often negative results from our experiments?
+The evaluation protocol of prior RGD-based analyses involved guiding generation with a RM, and then evaluating the resulting generations using the same RM Khanov et al.
+(2024).
+The issue with win-rates scored in this way is circularity.
+If RGD-adaptation can hack the RM (find a ’false positive’ response that the RM accepts, while not actually reflecting user preferences), using the same RM to evaluate the result produces overly optimistic results.
+To study this protocol, we report win Rate vs zero-shot in Table 5, which confirms the risk of ‘circular’ evaluation.
+Using the same RM to guide decoding and judge completions vs a baseline suffers from reward hacking/overoptimization El-Mhamdi and Hoang (2024).
+GenARM claims 100% improvement over zero-shot despite ROUGE-1 being marginally worse.
+Global RM claims 99.9% superiority while producing outputs identical to methods 20 points lower in RM accuracy.
+The circular evaluation protocol is thus vulnerable to reward hacking, and only appears to solve personal alignment.
+Despite their limitations, ranking held-out preferences (Table 2,4) does not suffer from this – because the RM is not used to generate; neither does our ground-truth evaluation (Figure.
+2) – because the RM generation is compared to ground-truth.
+Additional Analysis: ICL  We provide further analysis in Appendix C.1 showing ICL and ICL-RAG improve further with shots higher than 8 (as used in the main text).
 Scale RM Acc Policy Acc Win Rate
@@
 All correlations are negligible to negative across model scales, demonstrating no standard metric predicts behavioral alignment.
-Table 6 exposes the complete failure of standard evaluation: RM accuracy, policy accuracy, and win rates show negligible to negative correlations with generation quality (Kendall’s $\tau$ = -0.188 to -0.017).
-Even policy accuracy—measuring discrimination under the actual deployment mechanism—cannot predict behavioral alignment.
-Win rates are often most negative, confirming reward hacking.
-The takeaway is stark: no existing metric predicts whether methods will generate aligned outputs. Ground-truth evaluation on user-authored completions is not optional—it’s necessary.
-9 Discussion and Recommendations
-Our findings reveal a fundamental evaluation crisis: RM accuracy (Kendall’s $\tau$ = 0.08-0.31 with policy performance) and discrimination ability both fail to predict generation quality.
-Methods with 20-point RM accuracy gaps produce identical outputs, while reward models claiming 99% win rates show no ground-truth improvement.
-For practitioners: Use ICL-RAG for models $\geq$3B. Reward modeling adds complexity without benefit at scale.
-For researchers: (1) Report generation quality on ground-truth completions, not just RM accuracy.
-(2) Evaluate across model scales to detect scale-dependent failures.
-(3) Compare against ICL baselines.
-(4) Build behavioral benchmarks like Pref-LaMP with user-authored completions.
-The discrimination-generation gap reflects RGD’s nature as a complex dynamical system.
-At each generation step, two forces interact: base model logits and reward scores.
-These can agree or conflict, with dominance shifting across token positions.
-Policy accuracy measures whether the scoring function on average prefers tokens in chosen responses, but generation is sensitive to outliers—a single poorly-scored token can derail trajectories (causing length changes, topic drift).
-Moreover, discrimination and generation are distinct objectives: successfully ranking complete sequences doesn’t guarantee steering generation away from undesirable regions of the output space.
-Understanding these dynamics is critical future work.
-Limitations: We focus on inference-time methods (RGD) because it represents the only scalable deployment path—per-user RL fine-tuning remains computationally infeasible for realistic user populations.
-While RL-based personalization might exhibit different failure modes, it cannot address the fundamental problem we expose: practitioners need to select reward models for deployment, and current metrics fail to predict performance under the only viable deployment mechanism.
-Generalization beyond our three datasets requires investigation across diverse personalization dimensions (dialogue, values, humor).
-Our theoretical explanations of the discrimination-generation gap remain preliminary and warrant deeper mechanistic study.
-10 Conclusion
-We have demonstrated that personalized alignment from preference data faces a fundamental evaluation crisis: standard reward model accuracy metrics do not predict deployment performance under reward-guided decoding.
-Through systematic evaluation across three datasets and four model scales, we show that (1) RM accuracy correlates only weakly with discrimination ability, (2) discrimination ability doesn’t predict generation quality, and (3) simple in-context learning dominates reward-guided methods for models $\geq$3B parameters.
-These findings expose that the field has been optimizing proxy metrics divorced from actual deployment objectives.
-Current personalized alignment methods fundamentally fail to operationalize user preferences into behavioral adaptation under realistic deployment constraints.
-The introduction of Pref-LaMP—the first benchmark with ground-truth user completions—enables non-circular evaluation and reveals the extent of this failure through direct behavioral assessment.
-For practitioners, the path forward is clear: use ICL with retrieval for personalized generation at scale.
-For researchers, our work highlights the urgent need for ground-truth behavioral evaluation, theoretical understanding of the discrimination-generation gap, and development of methods that can effectively translate preferences into generation guidance.
-Only by measuring what actually matters—whether models generate outputs aligned with user preferences—can the field make progress on truly personalized alignment.
+No Metric Predicts Generation Quality  To summarise, we considered the standard RM accuracy, RGD-win rate, and our policy-accuracy metrics, all of which are discriminative ranking metrics.
+Table 6 correlates each of these against our end-to-end generation quality metric.
+All exhibit negligible to negative correlations with generation quality (Kendall’s $\tau$ = -0.188 to -0.017), with the negative correlations suggesting reward hacking.
+Takeaway: No existing metric predicts whether personalization methods will generate aligned outputs. Ground-truth evaluation on user-authored completions is a necessary evaluation component.
+9 Discussion and Conclusion
+Our findings reveal an evaluation crisis in personal alignment research: RM accuracy is uncorrelated with policy accuracy ($\tau$ = 0.08-0.31), and method rankings can completely invert between upstream and downstream evaluations.
+Using Pref-LaMP—the first benchmark with ground-truth user completions, we show discriminative metrics fail to predict generation quality (Table 6): reward models claiming 99% win rates show no improvement over baselines in ground-truth similarity.
+The field has been optimizing proxy metrics divorced from deployment objectives.
+On a more positive note, we highlight that in contrast to these issues with personal rewards and their evaluations, simple in-context learning dominates reward-guided methods for models $\geq$3B parameters, while being easy and reliable to implement.
+Practical Recommendations.
+Practitioners should use ICL with retrieval for 3B+ models; reward modeling adds complexity without benefit at scale.
+Researchers should: (1) evaluate complete pipelines end-to-end, not just reward model accuracy, (2) include policy accuracy and ground-truth behavioral metrics, (3) test across model scales to detect scale-dependent effects, (4) build behavioral benchmarks with user-authored completions and (5) compare against ICL baselines and focus future research effort on developing such amortized approaches to personal alignment.
+9.1 Limitations
+We focus on RGD because it represents a key scalable deployment path—per-user RL fine-tuning remains computationally infeasible for realistic populations.
+A fundamental challenge with RGD is that it assumes reward models can be token-wise factorized to provide local guidance at each generation step, which is a known source of error when this assumption is violated Li et al.
+(2025).
+While GenARM is specifically designed to address this limitation through token-level autoregressive reward training, it still exhibits the same performance gaps we observe across other methods.
+This suggests the problem runs deeper than factorization alone—the disconnect between preference learning and generation guidance may be fundamental to the inference-time adaptation paradigm.
+Our use of three datasets goes beyond most prior work, which often used only one or contrived datasets.
+However, our results do show some facets of dataset dependence, so future work should aim to establish larger multi-dataset benchmark suits to thoroughly test personalization across more dimensions of interest.
 Appendix A Extended Related Work and Limitations
@@
 Table 7: Reward model accuracy and policy accuracies per model and method for the PRISM dataset.
-LoRE LoRE-Alt MPU MPU-Avg PAL PReF VPL Global RM Global RM-v2
+LoRE LoRE-Alt MPU MPU-Avg PAL PReF VPL Global RM Global RM2
 RM Accuracy $56.53$ $\pm 0.15$ $60.33$ $\pm 0.57$ $49.66$ $\pm 1.13$ $49.07$ $\pm 1.46$ $62.51$ $\pm 0.30$ $62.73$ $\pm 0.64$ $60.31$ $\pm 2.98$ $64.51$ $\pm 0.41$ $59.18$ $\pm 0.74$
@@
 GenARM 83.89 $\pm$ 0.60 71.62 $\pm$ 0.85 40.82 $\pm$ 0.50 35.13 $\pm$ 0.42 68.93 $\pm$ 0.25
+Global RM 84.96 $\pm$ 0.13 66.41 $\pm$ 0.13 41.37 $\pm$ 0.02 35.87 $\pm$ 0.02 68.45 $\pm$ 0.00
+Global RM-v2 84.69 $\pm$ 0.07 79.13 $\pm$ 0.57 31.41 $\pm$ 0.75 26.40 $\pm$ 0.79 67.52 $\pm$ 0.56
 LoRE 65.60 $\pm$ 7.89 67.84 $\pm$ 0.50 41.40 $\pm$ 0.01 35.86 $\pm$ 0.01 68.45 $\pm$ 0.01
@@
 VPL 43.63 $\pm$ 3.77 63.58 $\pm$ 4.02 31.04 $\pm$ 14.71 26.68 $\pm$ 12.78 64.09 $\pm$ 5.90
-Vanilla RM 84.96 $\pm$ 0.13 66.41 $\pm$ 0.13 41.37 $\pm$ 0.02 35.87 $\pm$ 0.02 68.45 $\pm$ 0.00
-Vanilla RM-v2 84.69 $\pm$ 0.07 79.13 $\pm$ 0.57 31.41 $\pm$ 0.75 26.40 $\pm$ 0.79 67.52 $\pm$ 0.56
 DPO
@@
 Zero-Shot - - 40.32 $\pm$ 0.00 34.77 $\pm$ 0.00 67.70 $\pm$ 0.00
-ICL - nan $\pm$ nan 38.17 $\pm$ 0.20 33.52 $\pm$ 0.21 67.14 $\pm$ 0.13
+ICL - 69.46 $\pm$ 0.16 38.17 $\pm$ 0.20 33.52 $\pm$ 0.21 67.14 $\pm$ 0.13
 ICL-RAG - - 39.95 $\pm$ 0.02 35.62 $\pm$ 0.02 68.21 $\pm$ 0.00
@@
 GenARM 83.89 $\pm$ 0.60 71.24 $\pm$ 0.52 43.54 $\pm$ 0.74 37.53 $\pm$ 0.79 69.99 $\pm$ 0.04
+Global RM 84.96 $\pm$ 0.13 67.17 $\pm$ 0.04 45.47 $\pm$ 0.02 39.45 $\pm$ 0.03 70.32 $\pm$ 0.00
+Global RM-v2 84.69 $\pm$ 0.07 78.73 $\pm$ 0.71 32.75 $\pm$ 0.78 27.77 $\pm$ 0.78 68.27 $\pm$ 0.59
 LoRE 65.60 $\pm$ 7.89 68.49 $\pm$ 0.26 45.47 $\pm$ 0.03 39.45 $\pm$ 0.04 70.32 $\pm$ 0.01
@@
 VPL 43.63 $\pm$ 3.77 64.72 $\pm$ 3.68 35.19 $\pm$ 15.65 30.32 $\pm$ 13.94 66.08 $\pm$ 6.12
-Vanilla RM 84.96 $\pm$ 0.13 67.17 $\pm$ 0.04 45.47 $\pm$ 0.02 39.45 $\pm$ 0.03 70.32 $\pm$ 0.00
-Vanilla RM-v2 84.69 $\pm$ 0.07 78.73 $\pm$ 0.71 32.75 $\pm$ 0.78 27.77 $\pm$ 0.78 68.27 $\pm$ 0.59
 DPO
@@
 Zero-Shot - - 43.83 $\pm$ 0.00 37.89 $\pm$ 0.00 69.42 $\pm$ 0.00
-ICL - nan $\pm$ nan 43.09 $\pm$ 0.25 37.90 $\pm$ 0.07 69.50 $\pm$ 0.14
+ICL - 72.98 $\pm$ 0.38 43.09 $\pm$ 0.25 37.90 $\pm$ 0.07 69.50 $\pm$ 0.14
 ICL-RAG - - 43.96 $\pm$ 0.03 39.16 $\pm$ 0.01 69.96 $\pm$ 0.00
@@
 GenARM 83.89 $\pm$ 0.60 71.34 $\pm$ 0.56 43.30 $\pm$ 0.96 37.22 $\pm$ 0.99 69.94 $\pm$ 0.13
+Global RM 84.96 $\pm$ 0.13 68.28 $\pm$ 0.00 45.14 $\pm$ 0.03 39.20 $\pm$ 0.02 70.53 $\pm$ 0.01
+Global RM-v2 84.69 $\pm$ 0.07 78.02 $\pm$ 0.94 36.43 $\pm$ 1.99 30.36 $\pm$ 1.77 68.25 $\pm$ 0.43
 LoRE 65.60 $\pm$ 7.89 69.65 $\pm$ 0.31 45.14 $\pm$ 0.03 39.21 $\pm$ 0.02 70.53 $\pm$ 0.01
@@
 VPL 43.63 $\pm$ 3.77 66.75 $\pm$ 2.33 36.00 $\pm$ 14.93 30.90 $\pm$ 13.64 66.71 $\pm$ 6.07
-Vanilla RM 84.96 $\pm$ 0.13 68.28 $\pm$ 0.00 45.14 $\pm$ 0.03 39.20 $\pm$ 0.02 70.53 $\pm$ 0.01
-Vanilla RM-v2 84.69 $\pm$ 0.07 78.02 $\pm$ 0.94 36.43 $\pm$ 1.99 30.36 $\pm$ 1.77 68.25 $\pm$ 0.43
 DPO
@@
 Zero-Shot - - 44.86 $\pm$ 0.00 38.67 $\pm$ 0.00 70.28 $\pm$ 0.00
-ICL - nan $\pm$ nan 45.63 $\pm$ 0.04 40.14 $\pm$ 0.05 70.47 $\pm$ 0.06
+ICL - 75.67 $\pm$ 0.44 45.63 $\pm$ 0.04 40.14 $\pm$ 0.05 70.47 $\pm$ 0.06
 ICL-RAG - - 46.82 $\pm$ 0.03 41.38 $\pm$ 0.04 71.14 $\pm$ 0.00
@@
 GenARM 83.89 $\pm$ 0.60 72.08 $\pm$ 0.58 45.17 $\pm$ 0.53 39.18 $\pm$ 0.56 70.52 $\pm$ 0.11
+Global RM 84.96 $\pm$ 0.13 69.08 $\pm$ 0.04 45.96 $\pm$ 0.04 40.15 $\pm$ 0.02 70.75 $\pm$ 0.01
+Global RM-v2 84.69 $\pm$ 0.07 78.44 $\pm$ 0.90 38.91 $\pm$ 4.31 33.43 $\pm$ 3.62 68.96 $\pm$ 0.91
 LoRE 65.60 $\pm$ 7.89 70.97 $\pm$ 0.35 45.97 $\pm$ 0.03 40.16 $\pm$ 0.03 70.75 $\pm$ 0.00
@@
 VPL 43.63 $\pm$ 3.77 67.82 $\pm$ 2.50 37.57 $\pm$ 14.22 32.64 $\pm$ 12.70 67.53 $\pm$ 5.39
-Vanilla RM 84.96 $\pm$ 0.13 69.08 $\pm$ 0.04 45.96 $\pm$ 0.04 40.15 $\pm$ 0.02 70.75 $\pm$ 0.01
-Vanilla RM-v2 84.69 $\pm$ 0.07 78.44 $\pm$ 0.90 38.91 $\pm$ 4.31 33.43 $\pm$ 3.62 68.96 $\pm$ 0.91
 DPO
@@
 Zero-Shot - - 45.46 $\pm$ 0.00 39.23 $\pm$ 0.00 70.20 $\pm$ 0.00
-ICL - nan $\pm$ nan 47.42 $\pm$ 0.22 42.24 $\pm$ 0.26 71.54 $\pm$ 0.11
+ICL - 74.74 $\pm$ 0.25 47.42 $\pm$ 0.22 42.24 $\pm$ 0.26 71.54 $\pm$ 0.11
 ICL-RAG - - 48.90 $\pm$ 0.02 43.81 $\pm$ 0.01 72.20 $\pm$ 0.00
 Table 12: ROUGE scores for Qwen2.5-7B-Instruct (mean $\pm$ std across seeds).
+C.1 ICL Further Analysis
+We further analyse our strong ICL baselines in terms of number of demonstrations.
+ICL-RAG improves steadily with demonstrations and scale, reaching $\sim$49 ROUGE-1 at 7B with 8 shots.
+Larger models show no saturation, effectively leveraging context without reward guidance.
+This shows that personal alignment is not only possible, but straightforward to implement.
+However, operationalizing standard but more complex RM-based personal alignment approaches with RGD is comparatively fraught.
+This is shown in Figure 3.
+Figure 3: Pref-LAMP generation quality. ROUGE-1 vs. number of in-context examples. ICL-RAG consistently outperforms random ICL, with performance scaling with both model size and context length.
 Appendix D Architectural and Initialization Enhancements to PReF for Reward-Guided Decoding and Dataset Considerations
@@
                     response_2):
-    prompt1=prompts.PRISM_no_confidence.format(
+    prompt1=prompts.PRISM_no_confidence.\
+    format(
         prompt=prompt,
@@
 Our Choice: We prioritize authenticity by using real PRISM conversational preferences in their natural sparse form, tackling the more difficult—but more realistic—personalization problem.
+Appendix E LoRe Architecture
+LoRE is a pairwise preference learning method introduced in LoRe: Personalizing LLMs via Low-Rank Reward Modeling (lore2025).
+It learns a reward function from preference data, where each datapoint consists of a user input and two responses, one preferred over the other.
+Unlike methods that train a binary classifier to predict which response is better, LoRE optimizes a logistic loss over the difference of reward values assigned to the preferred and dispreferred responses.
+E.1 Architecture
+The LoRE architecture consists of two key components:
+Feature Extractor
+A shared feature extractor $\phi$ (typically a pretrained language model) processes the input $x$ and response $y$ to produce $K$ base reward scores: $R_{\phi}(x,y)\in\mathbb{R}^{K}$.
+User-Specific Weights
+For each user, we learn a low-rank weight vector $w\in\mathbb{R}^{K}$ that linearly combines these base rewards to produce a personalized scalar reward:
+$R_{w}(x,y)=w^{\top}R_{\phi}(x,y)$ (9)
+This architecture allows the model to learn a shared representation of reward dimensions through $\phi$, while capturing individual user preferences through the lightweight weight vectors $w$.
+E.2 Original Loss Formulation
+For a preference pair $(x,y^{+},y^{-})$ where $y^{+}$ is preferred over $y^{-}$, the loss uses the difference of personalized rewards:
+$\mathcal{L}_{\text{LoRE}}=\log(1+\exp(-w^{\top}[R_{\phi}(x,y^{+})-R_{\phi}(x,y^{-})]))$ (10)
+This encourages the model to assign a higher personalized reward to the preferred response $y^{+}$ over the dispreferred one $y^{-}$.
+E.3 Two Training Algorithms
+The paper introduces two variants:
+LoRE
+Trains both user-specific weights $w$ and the feature extractor $\phi$ simultaneously in a single optimization step. This approach was used for the TL;DR dataset in the original implementation.
+LoRE-Alt
+Uses an alternating optimization strategy: for each batch, it takes one gradient step on the user-specific weights $w$ (freezing the feature extractor $\phi$), then one step on the feature extractor $\phi$ (freezing the weights $w$).
+This approach was used for more complex datasets in the original implementation.
+LoRE-Alt also leverages an off-the-shelf reward model (Skywork RM) and includes a regularization term to prevent the learned model from deviating too far from the pretrained baseline.
+However, since we train our Qwen2.5-0.5B model from scratch without a pretrained reward model, we omit this regularization.
+Note: The original codebase does not successfully reproduce results on the PRISM dataset.
+E.4 our Variant: Equivalent Log-Sigmoid Loss
+In our implementation, we instead use:
+loss = -F.logsigmoid(reward_diff).mean()
+where:
+reward_diff = w.T @ (
+    R_phi(x, y^+) - R_phi(x, y^-)
+This is mathematically equivalent to the original logistic loss, since:
+$\log(\sigma(x))=\log(1+\exp(-x))$ (11)
+The logsigmoid loss is a numerically stable, PyTorch-friendly implementation of the same core principle.
+This change does not affect the training dynamics or final optimization target—it is purely an implementation detail.
+E.5 Architectural Equivalence
+In the original LoRE paper, the reward model can be formulated to take the difference of features directly:
+$w^{\top}[\phi(x,y^{+})-\phi(x,y^{-})]$ (12)
+In our implementation, we compute the reward separately on each response using the $K$-dimensional feature extractor, then take the weighted difference:
+$w^{\top}R_{\phi}(x,y^{+})-w^{\top}R_{\phi}(x,y^{-})$ (13)
+These formulations are mathematically identical due to the linearity of the inner product:
+$w^{\top}[R_{\phi}(x,y^{+})-R_{\phi}(x,y^{-})]=$ (14)
+$w^{\top}R_{\phi}(x,y^{+})-w^{\top}R_{\phi}(x,y^{-})$ (15)
+This equivalence holds because the personalization layer (the $w$ weights) is linear in the feature space.
+E.6 ARGS Support
+LoRE also supports Alignment as Reward-Guided Search (ARGS), where generation is guided at decoding time using the learned reward model.
+In our implementation, we enable ARGS as a runtime decoding strategy by plugging in the learned reward model as a plug-and-play scoring function.
+This is implemented by scoring candidate continuations during beam or sampling-based decoding using the personalized reward:
+$R_{w}(x,y_{\text{candidate}})=w^{\top}R_{\phi}(x,y_{\text{candidate}})$ (16)
+This allows us to steer generation toward responses that maximize the learned user-aligned reward signal, without requiring reinforcement learning or sampling from a reward-shaped distribution.
+E.7 Known Reproduction Issues
+It is a known issue that LoRe released code does not reproduce their results on the PRISM dataset due to issues in dataset preparation that confalted reported results.
+[^2] https://github.com/facebookresearch/LoRe/issues/1.
+This can be found here.
+Appendix F Hyperparameters
+F.1 Reward Modelling
+For the TLDR dataset, all models were trained with a LoRA of rank 8 and LoRA alpha of 16.
+rsLora was used for initializaiton.
+The backbone (LoRA module) was $5\times 10^{-5}$.
+Different models decoder used varying hyperparameters as listed below:
+  1. 1.
+    LoRe: decoder LR=0.5, latent dim=2
+  2. 2.
+[…]
+  3. 3.
+[…]
+  4. 4.
+    P-DPO: Number of soft tokens = 8
+  5. 5.
+    PReF: decoder LR was $1\times 10^{-3}$ and weight decay of 0.02. Latent dimension was 2
+  6. 6.
+[…]
+  7. 7.
+[…]
+For PRISM, and Pref-LaMP5 dataset, same hyperparameters were used with the differences of:
+  1. 1.
+    LoRe: latent dimension = 8
+  2. 2.
+    LoRe-Alt: same learning rates but latent dimension = 20
+  3. 3.
+    PReF: latent dimension = 8
+  4. 4.
+    PAL: latent dimension = 8
+F.2 ARGS
+To find the best weight hyperparameter for ARGS on TLDR and PRISM dataset, we maximized the policy accuracy metric over the seen/train users test split for each trained model.
+We found that this is necessary since reward models trained using the same algorithm can converge to producing rewards of different magnitudes.
+For LaMP-5, we used the harmonic mean of rouge-1, inverse perplexity (coherence), and policy accuracy to find the best weight which is subsequently used during generation.
+Appendix G Use of AI
+We used AI assistance in two capacities during this work:
+Code development: We used Cursor IDE with AI-assisted code completion during implementation. All AI-generated code suggestions were manually reviewed and verified before integration into the codebase.
+Writing assistance: We used large language models to help articulate technical concepts and improve clarity of exposition.
+The conceptual content, experimental design, results, and conclusions are entirely our own work.
+AI assistance was limited to rephrasing and refining presentation of ideas we specified.
+All scientific claims, experimental results, and intellectual contributions in this paper are the original work of the authors.